### v1

# Vector Quantization using Gaussian Variational Autoencoder

Abstract
Vector quantized variational autoencoder (VQ-VAE) is a discrete auto-encoder that compresses images into discrete tokens. It is difficult to train due to discretization. In this paper, we propose a simple yet effective technique, dubbed Gaussian Quant (GQ), that converts a Gaussian VAE with certain constraint into a VQ-VAE without training. GQ generates random Gaussian noise as a codebook and finds the closest noise to the posterior mean. Theoretically, we prove that when the logarithm of the codebook size exceeds the bits-back coding rate of the Gaussian VAE, a small quantization error is guaranteed. Practically, we propose a heuristic to train Gaussian VAE for effective GQ, named target divergence constraint (TDC). Empirically, we show that GQ outperforms previous VQ-VAEs, such as VQGAN, FSQ, LFQ, and BSQ, on both UNet and ViT architectures. Furthermore, TDC also improves upon previous Gaussian VAE discretization methods, such as TokenBridge. The source code is provided in https://github.com/tongdaxu/VQ-VAE-from-Gaussian-VAE.
1 Introduction
Vector-quantized variational autoencoder (van2017neural) is an autoencoder that compresses images into discrete tokens. It is fundamental to autoregressive generative models (esser2021taming; chang2022maskgit; yu2023language; sun2024autoregressive). However, VQ-VAE is difficult to train: the encoding process of VQ-VAE is not differentiable and challenges such as codebook collapse often emerge (sonderby2017continuous). Special techniques are required to ensure the convergence of VQ-VAE, such as commitment loss (van2017neural), expectation maximization (EM) (roy2018theory), Gumbel-Softmax (jang2016categorical; maddison2016concrete; sonderby2017continuous), and entropy loss (yu2023language; zhao2024image).
In this paper, we circumvent the challenge of training VQ-VAE by converting a Gaussian VAE with certain constraint into a VQ-VAE without any training. More specifically, we propose Gaussian Quant (GQ), a simple yet effective method for training-free conversion. The core idea is to generate a codebook of one-dimensional Gaussian noise and, for each dimension of the posterior, select the codebook entry that is closest to the posterior mean. Theoretically, we show that as the logarithm of the codebook size exceeds the bits-back coding bitrate (hinton1993keeping; townsend2019practical) of the Gaussian VAE, the resulting quantization error is small. In other words, GQ and the Gaussian VAE exhibit similar rate-distortion performance. This result serves as the theoretical foundation of GQ and provides a principled guideline for selecting codebook sizes.
Practically, we introduce the target divergence constraint (TDC) to train a Gaussian VAE for efficient conversion. TDC encourages the Gaussian VAE to achieve the same Kullback–Leibler (KL) divergence for each dimension, corresponding to the bits-back coding bitrate. Empirically, we demonstrate that GQ with Gaussian VAE trained by TDC, outperforms previous VQ-VAEs such as VQGAN, FSQ, LFQ, and BSQ (van2017neural; mentzer2023finite; yu2023language; zhao2024image) in terms of reconstruction quality, using both UNet and ViT backbones. Additionally, we show that TDC can improve previous Gaussian VAE discretization methods, such as TokenBridge (wang2025bridging).
Figure 1: The rate-distortion performance on the ImageNet dataset demonstrates that GQ outperforms previous VQ-VAEs on both UNet and ViT architectures.
Our contributions can be summarized as follows:
  • (Section 3.1) We propose GQ, a simple yet effective approach that converts a pre-trained Gaussian VAE with certain constraint into VQ-VAE without training.
  • (Section 3.2) Theoretically, we prove that when the codebook size of GQ is close to the bits-back coding bitrate of the Gaussian VAE, the conversion error remains small.
  • (Section 3.3) Empirically, we introduce target divergence constraint (TDC) to implement GQ and show that GQ outperforms previous VQ-VAEs, such as VQGAN, FSQ, LFQ, and BSQ, on both UNet and ViT architectures.
  • (Section 3.5) Furthermore, we show that TDC can be used to improve previous Gaussian VAE discretization approaches, such as TokenBridge.
2 Preliminaries
We denote the original image as $X$, the latent variable as $Z$, the encoder as $f(\cdot)$, and the decoder as $g(\cdot)$. We use $\log(\cdot)$ to denote the natural logarithm (base $e$) and KL divergence as $D_{KL}(\cdot||\cdot)$. Similarly, we use $\log_{2}(\cdot)$ to denote the logarithm with base $2$, and KL divergence as $D_{KL(2)}(\cdot||\cdot)$.
2.1 Vector Quantized Variational Autoencoder
VQ-VAE (van2017neural) transforms source image into a series of integer tokens, which can be decoded using a codebook and decoder. To facilitate auto-regressive generation, it typically involves a deterministic transformation and a shared codebook across different tokens. More specifically, VQ-VAE maintains a codebook $c_{1:K}$ with size $K$ and a bitrate of $\log K$. The encoding process of VQ-VAE involves finding the closest codeword $c_{j}$ in $c_{1:K}$ to the encoder output $f(x)_{i}$ for each latent dimension $i$. Denote distortion as $\Delta(\cdot,\cdot)$, the optimization target of VQ-VAE is the rate-distortion function weighted by Lagrangian multiplier $\lambda$:
$\mathcal{L}_{VQ}=\lambda\underbrace{\log K}_{\textrm{bitrate}}+\underbrace{\mathbb{E}[\Delta(X,g(\hat{z}))]}_{\textrm{distortion}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{Reg}},$
[…]
and $\mathcal{L}_{Reg}$ are regularization for VQ-VAE to converge, such as a combination of commitment loss and codebook loss (van2017neural) and Gumbel Softmax loss (sonderby2017continuous).
2.2 Gaussian Variational Autoencoder and Bits-Back Coding
The Gaussian VAE is a special type of VAE (kingma2013auto) with a prior $\mathcal{N}(0,I)$ and a fully factorized Gaussian posterior $q(Z|X)$. The encoding process of a Gaussian VAE simply involves sampling the latent variable $z_{i}\sim q(Z_{i}|X)$ for each latent dimension $i$. Assuming $\log p(X|Z=z)\propto(1/\lambda)\Delta(X,g(z))$, then the negative evidence lower bound (ELBO) of the Gaussian VAE is equivalent to a rate-distortion function of a bits-back coding bitrate term and a distortion term:
$\mathcal{L}_{VAE}=\lambda\underbrace{D_{KL}(q(Z|X)||\mathcal{N}(0,1))}_{\textrm{bits-back coding bitrate}}+\underbrace{\mathbb{E}[\Delta(X,g(z))]}_{\textrm{distortion}},$
$z_{i}\sim q(Z_{i}|X=x)=\mathcal{N}(\mu_{i},\sigma_{i}^{2}),i=1\ldots d.$ (2)
The bitrate of $z_{i}$ is the bits-back coding bitrate, defined as $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,I))$ (hinton1993keeping; townsend2019practical). This is because, when compressing $X$ losslessly, one can communicate $z_{i}$ using $D_{KL}(q(Z|X)||\mathcal{N}(0,I))$ nats for arbitrary precision.
3 Gaussian Quant: Vector Quantization using Gaussian VAE
3.1 Direct Quantization of Gaussian VAE
We propose an extremely simple technique to obtain a VQ-VAE from a Gaussian VAE: we directly generate one-dimensional Gaussian noise as the codebook for VQ-VAE (van2017neural) and quantize the posterior mean $\mu_{i}$ of the Gaussian VAE independently for each dimension $i$. Because the codebook consists entirely of samples from a Gaussian distribution, we refer to our approach as Gaussian Quant (GQ). Specifically, we randomly generate $K$ codebook values $c_{1:K}\sim\mathcal{N}(0,1)$, which is the same for each dimension. Then, for each dimension $i$, we select the $c_{j}$ that is closest to the posterior mean $\mu_{i}$ and denote the quantized value as $\hat{z}_{i}$:
[…]
3.2 Theoretical Relationship between the Codebook Size and Quantization Error
Why GQ works and how to select $K$ are not straightforward questions. Theoretically, we show that GQ preserves the rate-distortion property of the Gaussian VAE: when the bitrate $\log K$ matches the bits-back coding bitrate of the Gaussian VAE, the quantization error is small. More specifically, we show that the probability of large quantization error decays doubly exponentially as the codebook bitrate $\log K$ exceeds the bits-back coding bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$.
Theorem 1.
Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i}$ and $\sigma_{i}$, respectively. Assuming that the product and sum satisfy $|\mu_{i}\sigma_{i}|\leq c_{1}$ and $|\mu_{i}|+|\sigma_{i}|\leq c_{2}$, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays doubly exponentially with respect to the number of nats $t$ by which the codebook bitrate $\log K$ exceeds the bits-back coding bitrate. i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-e^{t}\underbrace{\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}}_{\textup{constant}})}.$ (4)
Conversely, when the codebook bitrate $\log K$ is smaller than the bits-back coding bitrate, the probability of large quantization error increases exponentially toward $1$. More specifically, we show that the probability of large quantization error increases exponentially when the codebook bitrate $\log K$ is lower than the bits-back bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$.
Theorem 2.
The probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increases exponentially with respect to the number of nats $t$ by which the codebook bitrate $\log K$ is lower than the bits-back coding bitrate, i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\geq 1-e^{-t}\underbrace{\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}}_{\textup{constant}}.$ (5)
Theorems 1 and 2 provide a principled guideline for choosing $K$, such that $\log K$ should be close to the bits-back bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$. In practice, setting $\log_{2}K=\left\lceil D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\right\rfloor$ typically yields a small enough reconstruction error, where $\left\lceil\cdot\right\rfloor$ denotes the rounding operator. Using a larger $K$ does not provide additional benefits, while a smaller $K$ increases the error significantly.
3.3 Practical Implementation with Target Divergence Constraint
There are two challenges to make GQ practical. The first challenge is, if we want to construct a VQ-VAE with a specific codebook size $K$, how can we train a Gaussian VAE with corresponding KL divergence? The second challenge is, for a vanilla Gaussian VAE trained to minimize the loss in Eq.2, the values of $D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))$ vary significantly across dimensions $i$. How to train a Gaussian VAE with KL divergence remains close to $\log K$ across each dimension?
[…]
$\mathcal{L}_{TDC}=\sum_{i=1}^{d}\lambda_{i}D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+\Delta(X,g(z)),$
$\textup{where }\lambda_{i}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textup{ bits},\\ \lambda_{\textrm{mean}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\in[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha,\log_{2}K+\alpha}]\textup{ bits},\\ \lambda_{\textrm{max}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textup{ bits}.\end{array}\right$. (9)
To determine $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$, we extend the heuristic in MIRACLE (havasi2018minimal) and HiFiC (Mentzer2020HighFidelityGI). More specifically, we initialize $\lambda_{\textrm{min}}=\lambda_{\textrm{mean}}=\lambda_{\textrm{max}}=1$, and update them according to the following rule:
$\lambda_{\textrm{min}}=\lambda_{\textrm{min}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ min}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textrm{ else }\lambda_{\textrm{min}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{mean}}=\lambda_{\textrm{mean}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ mean}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>\log_{2}K\textrm{ else }\lambda_{\textrm{mean}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{max}}=\lambda_{\textrm{max}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ max}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textrm{ else }\lambda_{\textrm{max}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$ (10)
where $\beta$ is the hyper-parameter controlling update speed. To avoid numerical issue, we further clip $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$ into a range of $10^{-3},10^{3}]$ after each update. In practice, we use $\alpha=0.5,\beta=1.01$. In Appendix B, we propose an alternative implementation of TDC using Lambert W function, which is less effective for ViT models.
3.4 Grouping to Multiple Dimensions
The vanilla VQ-VAE has three key parameters: codebook size, codebook dimension and number of token. Some previous alternative to VQ-VAE, such as LFQ and BSQ (yu2023language; zhao2024image) limit the codebook dimension to 1. To support codebook dimension greater than 11, we can group $m$ tokens into a single large token with codebook size $\log_{2}K=\left\lceil\sum_{l=i}^{i+m}D_{KL}(q(Z_{l}|X)||\mathcal{N}(0,1))\right\rfloor$. There are three grouping strategies to achieve this: post-quantization (PQ), post-training (PT), and training-based (TR), with different trade-offs in flexibility and performance. We brief these strategies in main text and provide details in Appendix C.2.
PQ is the most flexible grouping strategy and can be applied after GQ. For PQ, since the posterior of the Gaussian VAE $q(Z|X)$ is a factorized Gaussian, the quantization is also independent across dimensions. This means that we can trivially combine $m$ tokens into a larger one, by treating each token as an integer in $K^{1/m}$-based number system and aggregate them, which is the same as other one dimensional VQ-VAEs (chang2022maskgit; mentzer2023finite; zhao2024image).
PT is less flexible than PQ, as it can only be applied before GQ and after training the Gaussian VAE. For PT, we can view the one-dimensional GQ in Eq. 3 as the maximum likelihood estimator of a one-dimensional Gaussian. This approach can be extended to an $m$-dimensional diagonal Gaussian distribution. Additionally, for low bitrate cases, we observe that $m$-dimensional PT leads to low codebook usage. This is because $|\mu_{i}|$ is bounded by $\sqrt{2D_{KL}}$. When the $D_{KL}$ is small, some vector in the codebook that is far from 0 is never used. To address this, we introduce a regularization term weighted by $\omega$ to improve codebook usage by encouraging the selection $c_{j}$ that is far from 0:
[…]
TR is the least flexible strategy and must be used during the training of the Gaussian VAE. Specifically, building on the PT quantization approach, we can relax the TDC training target by considering the relationship between $\sum_{l=i}^{i+m}D_{KL}(q(Z_{l}|X)||\mathcal{N}(0,1))$ and $\log_{2}K\pm 0.5$ bits. In terms of performance, PQ does not affect reconstruction at all. PT provides slight improvements in reconstruction, while TR significantly enhances reconstruction performance (see Table 13).
3.5 Improving TokenBridge with Target Divergence Constraint
TokenBridge (wang2025bridging) also convert a pre-trained Gaussian VAE into a VQ-VAE. It adopts the Post Training Quantization (PTQ) technique from model compression and proposes to treat latent as model parameters to discretize. It uses a fixed codebook composed of $2^{K}$ centroids of a Gaussian distribution. It then quantizes the posterior sample by finding the closest centroid. However, TokenBridge directly quantizes a vanilla Gaussian VAE without limiting the KL divergence of each dimension. This leads to suboptimal rate-distortion performance. In fact, we can also improve the performance of TokenBridge using TDC. Specifically, the quantization centers of TokenBridge are the equal-probability partition centers of the $\mathcal{N}(0,1)$ distribution, which can be seen as a special case of GQ with an evenly distributed codebook $c_{1:K}$. From this perspective, the number of PTQ bits should also match the bits-back coding bitrate of the Gaussian VAE, and TDC can therefore enhance the performance of TokenBridge. As we demonstrate in Table 3, TDC indeed improves TokenBridge performance by a large margin.
3.6 Relationship with Reverse Channel Coding
GQ is closely related to reverse channel coding, which aims to simulate a distribution $q$ using samples from a distribution $p$ (harsha2007communication; li2018strong; havasi2018minimal; Flamich2020CompressingIB; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating). For example, Minimal Random Coding (MRC) (havasi2018minimal), when applied to a Gaussian VAE, samples from the categorical distribution with logits given by the likelihood difference:
$\hat{z}_{i}\sim\hat{q}(c_{1:K})\propto e^{\log q(Z_{i}=c_{j}|X)-\log\mathcal{N}(c_{j}|0,1)}.$ (12)
The key difference between MRC and GQ is that MRC and its variants (havasi2018minimal; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating) simulate a distribution through stochastic sampling, whereas a VQ-VAE requires deterministic quantization. For one dimension quantization, the bias bound of MRC derived from Chatterjee2015TheSS can not be achieved as VQ-VAE does not allow stochastic encoding. On the other hand, our achievability and converse bound is compatible with deterministic quantization. In terms of quantization error, GQ outperforms MRC by definition (Eq. 3). Besides, GQ without grouping (m=1) can be implemented by bisect search with better asymptotic complexity (See Appendix D.11).
Additionally, TDC is closely related to the MIRACLE heuristic and IsoKL parametrization of Gaussian VAE (Havasi2018MinimalRC; Flamich2022FastRE; Lin2023MinimalRC). More specifically, MIRACLE also proposes adjusting $\lambda$ during VAE training. However, MIRACLE only maintains a single $\lambda$, making it less effective for controlling the mean of $D_{KL}$ but less effective for constraining its minimum and maximum values. On the other hand, IsoKL imposes strict control on $D_{KL}$ by directly solving for $\sigma$ given $\mu$ using the Lambert W function (corless1996lambert; brezinski1996extrapolation). However, IsoKL suffers from numerical issues and leads to suboptimal performance.
Figure 2: Qualitative results on ImageNet dataset and 0.25 bpp. Our GQ has most visually pleasing reconstruction result.
4 Experimental Results
4.1 Experimental Setup
Models and Baselines For image reconstruction, we select two representative autoencoder architectures: UNet from Stable Diffusion 3 (Esser2024ScalingRF), and ViT from the BSQ (zhao2024image). For VQ-VAE baselines, we include vanilla VQGAN (van2017neural) and several representative variants, including FSQ (mentzer2023finite), LFQ (yu2023language), and BSQ (zhao2024image). Besides, we compare our approach to pre-trained VQ-VAE of VQGAN-Taming (esser2021taming), VQGAN-SD (rombach2022high) Llama-Gen (sun2024autoregressive), FlowMo (Sargent2025FlowTT), BSQ (zhao2024image) and other conversion approaches such as TokenBridge and ReVQ . Additionally, we demonstrate that our TDC technique can improve TokenBridge (wang2025bridging), a previous training-free approach for converting Gaussian VAEs into VQ-VAE. For image generation, we employ the Llama transformer (Touvron2023Llama2O; Shi2024ScalableIT).
Datasets, Bitrates, and Metrics For the datasets, we use the ImageNet (Deng2009ImageNetAL) training split for training, and both the ImageNet and COCO (Lin2014MicrosoftCC) validation split for testing. For reconstruction and generation experiments, all images are resized to $256\times 256$ and $128\times 128$ respectively. In terms of bitrates, we evaluate the image reconstruction performance of all models using a codebook size of $2^{14}-2^{18}$ and token numbers of $1024$, $2048$, and $4096$, which correspond to bpp (bits-per-pixel) values of 0.221.000.22-1.00. This extends the BSQ evaluation beyond 0.250.500.25-0.50 bpp. For metrics, we adopt Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang2018TheUE), Structural Similarity Index Measure (SSIM) (Wang2004ImageQA), and reconstruction Fréchet Inception Distance (rFID) (Heusel2017GANsTB) for image reconstruction; generation Fréchet Inception Distance (gFID) and Inception Score (IS) (Salimans2016ImprovedTF) are used for image generation. For further details, see Appendix C.
Table 1: Quantitive results on ImageNet dataset. Our GQ outperforms other VQ-VAEs on both UNet and ViT architecture, across 0.25-1.00 bpp. Bold: best.
Method
bpp
(# of tokens)
UNet based ViT based
[…]
VQGAN
0.25
[…]
26.51 0.125 0.748 5.714 25.39 0.103 0.740 3.518
FSQ 26.34 0.075 0.756 1.125 25.87 0.109 0.751 3.856
LFQ 24.74 0.164 0.722 16.337 24.81 0.143 0.725 15.716
BSQ 25.62 0.086 0.754 1.080 26.52 0.083 0.793 1.649
GQ (Ours) 27.61 0.059 0.807 0.529 27.88 0.061 0.823 0.932
VQGAN
0.50
($^{\textup{16}}\times2$048)
29.21 0.052 0.831 1.600 27.86 0.062 0.823 1.228
FSQ 29.29 0.047 0.845 0.871 28.83 0.055 0.842 1.067
LFQ 26.90 0.107 0.800 8.035 27.87 0.068 0.829 2.444
BSQ 27.88 0.059 0.836 0.788 28.44 0.051 0.852 0.700
GQ (Ours) 30.17 0.039 0.875 0.492 30.42 0.037 0.882 0.592
VQGAN
1.00
[…]
32.06 0.026 0.896 0.580 31.32 0.032 0.899 0.716
FSQ 32.38 0.025 0.905 0.636 31.58 0.026 0.905 0.544
LFQ 28.31 0.074 0.840 3.617 26.67 0.105 0.790 8.288
BSQ 30.50 0.032 0.900 0.346 31.60 0.027 0.914 0.379
GQ (Ours) 32.47 0.023 0.907 0.322 31.71 0.024 0.903 0.349
Table 2: Quantitive results on ImageNet dataset. Our GQ outperforms previous pre-trained VQ-VAEs with less training. Bold: best, : from paper, -: not available.
Method
bpp
(# of tokens)
[…]
ImageNet
[…]
VQGAN-Taming
0.22
[…]
23.38 - - 1.190 (OpenImages) 67
VQGAN-SD - - - 1.140 (OpenImages) 83
Llama-Gen-32 24.44 0.064 0.768 0.590 40 70
FlowMo-Hi 24.93 0.073 0.785 0.560 300 945
GQ (Ours) 25.31 0.064 0.762 0.491 40 82
BSQ
0.28
[…]
27.78 0.063 0.817 0.990 200 175
GQ (Ours) 27.86 0.054 0.804 0.424 40 87
Table 3: Quantitive results of improving TokenBridge based on UNet architecture. Our TDC improves TokenBridge significantly.
Method
bpp
(# of tokens)
ImageNet validation COCO validation
[…]
Gaussian VAE
$\approx$ 1.00 (-)
32.73 0.022 0.910 0.490 32.64 0.018 0.917 2.380
Gaussian VAE (w/ TDC) 32.61 0.023 0.906 0.460 32.69 0.019 0.919 2.717
TokenBridge
1.00
($2^{16}\times 4096$)
28.24 0.045 0.869 0.823 28.19 0.043 0.878 4.167
TokenBridge (w/ TDC) 31.67 0.025 0.903 0.385 31.56 0.022 0.910 2.171
GQ (Ours) 32.60 0.022 0.908 0.280 32.53 0.020 0.917 2.153
4.2 Main Results
Image Reconstruction In Table 1 and Table 15, we compare our GQ method to other quantization approaches across the 0.250.251.001.00 bpp range. The results show that, in terms of reconstruction metrics such as PSNR, LPIPS, SSIM, and rFID, our GQ approach achieves state-of-the-art performance in most cases. The advantage of GQ is consistent across both UNet and ViT model architectures, as well as for both the ImageNet and COCO datasets. Visually, in Figure 2, it is shown that our GQ also produces pleasing reconstruction and preserves a lot more details in the source image. Besides, in Table. 2, we show that our GQ achieves competitive performance compared with several pre-trained models such as FlowMo, with less training epochs. Additionally, in Table. 11, we show that our GQ outperforms previous methods to discretize Gaussian VAE, including TokenBridge and ReVQ.
Improving TokenBridge In Table 3, we compare TokenBridge (wang2025bridging) applied to a vanilla Gaussian VAE and a TDC constrainted Gaussian VAE. The results show that the quantization error of TokenBridge is quite large when applied to a vanilla Gaussian VAE. In contrast, TDC significantly reduces the quantization error.
Image Generation In Table 4, we evaluate the performance of GQ in terms of image generation. It is shown that compared with VQGAN, FSQ, LFQ and BSQ, our GQ has higher codebook usage and codebook entropy. In terms of generation FID and IS, our GQ is comparable to FSQ and better than other methods. Additionally, we train a DiT (Peebles2022ScalableDM) with same model architecture and same training setting, using the Gaussian VAE with and without TDC. It is shown that for limited computation, auto-regressive generation is more efficient than diffusion generation, in terms of both FID and IS. This result shows that the conversion from Gaussian VAE to VQ-VAE facilities auto-regressive generation, which improves the efficiency of image generation.
Complexity Compared with Gaussian VAE, the overhead of GQ is negligible (See Appendix D.10).
Table 4: Quantitive results on class conditional image generation of ImageNet dataset. Our GQ achieves best codebook usage and competitive generation performance. Bold: best.
[…]
Diffusion
Gaussian VAE w/o TDC - - 8.35 202.19
Gaussian VAE w/ TDC - - 8.47 205.94
Auto-regressive
VQGAN 16.4% 4.36 8.01 151.40
FSQ 94.3% 14.74 7.33 224.88
LFQ 24.9% 9.65 7.73 142.09
BSQ 99.8% 14.93 7.82 221.64
TokenBridge 94.6% 14.94 7.82 198.24
GQ (Ours) 100.0% 15.17 7.67 230.79
4.3 Ablation Studies
Effectiveness of Pre-trained Gaussian VAE It is possible to train a vanilla VQ-VAE (van2017neural) using the same codebook as GQ, which is equivalent to a VQ-VAE with a fixed, Gaussian noise codebook. It is also possible to directly train the Gaussian VAE neural network with the GQ target in Eq. 11 using Gumbel-Softmax (jang2016categorical; maddison2016concrete). However, as shown in Table 5, both methods do not converge well. Furthermore, fine-tuning GQ after initializing it with a pre-trained Gaussian VAE also has only marginal effect on performance. These results indicate that GQ’s conversion from a pre-trained Gaussian VAE is necessary and sufficient.
Table 5: The effect of pre-trained Gaussian VAE. Converting GQ from pre-trained Gaussian VAE is better than training GQ from scratch.
[…]
GQ (from scratch) Vanilla VQ-VAE 1.00 8.50 0.763 0.156 360.597
GQ (from scratch) Gumbel-softmax Eq. 11 29.65 0.044 0.866 0.928
GQ (finetune from Gaussian VAE) Gumbel-softmax Eq. 11 32.45 0.022 0.905 0.264
GQ (convert from Gaussian VAE) no 32.47 0.023 0.907 0.327
Effectiveness and Alternatives of TDC To demonstrate the necessity of TDC in Eq. 9, we train a vanilla Gaussian VAE without TDC. As shown in Table 6, the mean $D_{KL(2)}$ of the vanilla Gaussian VAE is close to that of the Gaussian VAE with TDC (3.99 vs. 4.26 bits). However, the range of $D_{KL(2)}$ is much wider for the vanilla model (0.26–27.29 vs. 2.93–5.63 bits). Although the reconstruction performance of the two Gaussian VAEs is very similar (PSNR: 32.73 vs. 32.61 dB, rFID: 0.490 vs. 0.460), GQ with TDC outperforms GQ without TDC by a large margin (PSNR: 31.25 vs. 26.43 dB, rFID: 0.372 vs. 0.978). This demonstrates that TDC is necessary.
Alternatives to TDC are the MIRACLE heuristic (Havasi2018MinimalRC) and IsoKL (Flamich2022FastRE). MIRACLE is not that effective in terms of controlling the range is outperformed by TDC (PSNR 29.48 vs. 32.11 dB). On the other hand, IsoKL imposes a stricter constraint by requiring that $D_{KL}$ is exactly the same across all dimensions. Iso-KL enforece the constraint well but has inferior performance (PSNR 30.45 vs. 32.11 dB). This is because IsoKL is not numerically stable and it discards the solution with $\sigma^{2}_{i}>1$. In Appendix B, we propose a numerically stable version of Mean-KL (Lin2023MinimalRC), which is a IsoKL that supports grouping ($m>1$). However, it does not work well for ViT based models.
Table 6: The effect of adding constraints to Gaussian VAE. GQ is effective only on Gaussian VAE trained with TDC constraint.
[…]
Gaussian VAE no 3.99, 0.26-27.29 - 1.00 32.73 0.022 0.910 0.490
GQ - - 4 1.00 26.43 0.834 0.054 0.978
Gaussian VAE MIRACLE 4.34,0.91-26.98 - 1.00 32.82 0.023 0.910 0.436
GQ - - 4 1.00 29.48 0.039 0.887 0.439
Gaussian VAE IsoKL 4.34, 4.24-4.38 - 1.00 30.54 0.878 0.027 0.400
GQ - - 4 1.00 30.45 0.878 0.030 0.468
Gaussian VAE TDC (ours) 4.26, 2.93-5.63 - 1.06 32.61 0.023 0.906 0.460
GQ - - 4 1.00 32.11 0.023 0.906 0.414
Alternatives to GQ There are several stochastic alternatives to GQ, such as MRC, ORC, and A coding (havasi2018minimal; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating). In Table 7, we compare these methods in terms of reconstruction quality. When applied to a TDC-constrained Gaussian VAE, GQ achieves the best PSNR, SSIM, and rFID. Besides, when grouping $m=1$, GQ can be implemented using bisect search. This means that GQ is asymptotically faster than MRC, ORC and A coding (See Appendix D.11).
Table 7: Comparison between MRC methods and GQ on ImageNet dataset. GQ has better reconstruction quality and can be implemented asymptotically faster when grouping $m=1$.
[…]
Gaussian VAE (w/ TDC) O(1)O(1)/$O(1)$ 32.75 0.023 0.906 0.460
MRC (original) $O(2^{DKL(q(Z_{i}|X)||\mathcal{N}(0,1))})$/$O(1)$ 32.09 0.023 0.906 0.425
MRC (A coding) $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$$O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ 32.09 0.023 0.906 0.425
ORC $O(2^{DKL(q(Z_{i}|X)||\mathcal{N}(0,1))})$/$O(1)$ 32.09 0.023 0.906 0.419
GQ (Ours) $O(D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)))$/$O(1)$ 32.11 0.023 0.907 0.414
The TDC Parameters and Grouping Strategies See Appendix D.
5 Related works
VQ-VAE (van2017neural) is an autoencoder that compresses images into discrete tokens. Due to the discretization, it is impossible to be train directly using gradient descent. Various techniques have been proposed to address this, such as commitment loss (van2017neural), expectation maximization (EM) (roy2018theory), the straight-through estimator (STE) (bengio2013estimating), and Gumbel-softmax (jang2016categorical; maddison2016concrete; sonderby2017continuous; Shi2024TamingSV). In addition, VQ-VAE is prone to codebook collapse. To mitigate this, various methods have been proposed, such as reducing the code dimension (Yu2021VectorquantizedIM; Sun2024AutoregressiveMB), product quantization (Zheng2022MoVQMQ), residual quantization (Lee2022AutoregressiveIG), dynamic quantization (Huang2023TowardsAI), multi-level quantization (Razavi2019GeneratingDH), feature projection (Zhu2024ScalingTC), rotation codebooks (Fifty2024RestructuringVQ) and etc. (yu2021vector; chiu2022self; takida2022sq; zhang2023regularized; huh2023straightening; gautam2023soft; goswami2024hypervq).
More related to our work, some variants of VQ-VAE with fixed codebook emerge, such as FSQ (mentzer2023finite), LFQ (yu2023language), and BSQ (zhao2024image). However, training tricks such as the straight-through estimator (STE) is still required. Among all previous works, TokenBridge (wang2025bridging) and ReVQ (zhang2025quantizethenrectifyefficientvqvaetraining) are most aligned with our objective. TokenBridge and ReVQ also convert a pre-trained Gaussian VAE into a VQ-VAE. However, TokenBridge and ReVQ do not constraint the divergence of Gaussian VAE, leading to suboptimal performance. Besides, ReVQ requires some training while our approach is training-free.
Reverse Channel Coding See Section. 3.6.
6 Conclusion & Discussion
In this paper, we propose Gaussian Quant (GQ), an extremely simple yet effective technique that converts a pre-trained Gaussian VAE into a VQ-VAE without any additional training. Theoretically, we show that when the logarithm of the GQ codebook size exceeds the bits-back coding bitrate of the Gaussian VAE, a small quantization error is achieved. In addition, we propose a target divergence constraint (TDC) to implement GQ effectively. Empirically, we demonstrate that GQ outperforms previous discrete VAEs, such as VQGAN, FSQ, LFQ, and BSQ (van2017neural; mentzer2023finite; yu2023language; zhao2024image). Furthermore, our TDC also improves the performance of TokenBridge (wang2025bridging).
We limit our evaluation of GQ to the standard SD3.5 UNet (Esser2024ScalingRF) and the BSQ-ViT architecture. Additionally, we restrict the bpp range to 0.220.221.001.00, which extends the BSQ’s original range of 0.250.250.500.50 bpp. We acknowledge that there are several highly competitive VQ-VAEs that adopt multi-scale or residual architectures (Tian2024VisualAM; Han2024InfinitySB) or study the low bpp regime (bpp $\leq 0.2$) (Yu2024AnII; Sargent2025FlowTT; zhang2025quantizethenrectifyefficientvqvaetraining). However, in this paper, we use standard architectures and a typical bpp range to focus on the core aspects of the quantization method. Additionally, we focus on achieving a good trade-off between bitrate and reconstruction quality, leaving the complex relationship between reconstruction and generation performance to future works (wang2024image; xiong2025gigatok; hansen2025learnings).
Ethics Statement
The approach proposed in this paper focus on reconstruction of existing images with limited bitrate. As the model is essentially not generative, the ethic concerns is not obvious. Nevertheless, the GAN module in decoder might has negative effects, including issues related to mis-representation and trustworthiness.
Reproducibility Statement
For theoretical results, the proofs for all theorems are presented in Appendix A. For the experiments, we use publicly accessible datasets such as ImageNet (Deng2009ImageNetAL). Implementation details and hyper-parameters are provided in Appendix C. Besides, we include the source code for reproducing the experimental results in the supplementary material.
Appendix A Proof of Main Results
Theorem 1. Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i},\sigma_{i}$, and assuming $|\mu_{i}|,\sigma_{i}$ ’s product and sum are bounded by $|\mu_{i}\sigma_{j}|\leq c_{1}$, $|\mu_{i}|+|\sigma_{j}|\leq c_{2}$, then the probability of quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays double exponentially to the amount of nats $t$ that codebook bitrate $\log K$ exceeds the bits-back coding bitrate, i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-e^{t}\underbrace{\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}}_{\textup{constant}})}.$ (6)
Proof.
Denote the cumulative distribution function (CDF) of $\mathcal{N}(0,1)$ as $\Phi$, and probability density function (PDF) of $\mathcal{N}(0,1)$ as $\phi$, then we need to consider the probability that no samples falls between $\mu_{i}-\sigma_{i},\mu_{i}+\sigma_{i}]$, which is
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}=(1-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})))^{K}.$ (13)
Now we use the Bernoulli inequality, that $\forall y\in\mathbb{R},1+y\leq e^{y}$. Let $y=-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i}))$, we have
$1-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i}))\leq\exp{(-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})))}.$ (14)
Taking Eq. 14 into Eq. 13, we have
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})))}^{K}$ (15)
$=\exp{(-K\cdot(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})))}$
$=\exp{(-K\cdot\int_{\mu_{i}-\sigma_{i}}^{\mu_{i}+\sigma_{i}}\phi(x)dx)}.$
By integral mean value theorem, $\exists x^{\prime}\in[\mu_{i}-\sigma_{i},\mu_{i}+\sigma_{i}]$, such that
$\int_{\mu_{i}-\sigma_{i}}^{\mu_{i}+\sigma_{i}}\phi(x)dx=2\sigma_{i}\phi(x^{\prime}).$ (16)
And then we have
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\leq\exp{(-K\cdot 2\sigma_{i}\phi(x^{\prime}))}.$ (17)
Next, we consider three cases: $\mu_{i}-\sigma_{i}\geq 0$, $\mu_{i}+\sigma_{i}\leq 0$, and $\mu_{i}-\sigma_{i}\leq 0\leq\mu_{i}+\sigma_{i}$.
First, consider the case when $\mu_{i}-\sigma_{i}\geq 0$. Obviously we have $\phi(\mu_{i}+\sigma_{i})\leq\phi(x^{\prime})$, and we have
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot 2\sigma_{i}\phi(\mu_{i}+\sigma_{i}))}$
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}\sigma_{i}e^{-\frac{1}{2}(\mu_{i}+\sigma_{i})^{2}})}$
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(\mu_{i}^{2}+\sigma_{i}^{2}-\log\sigma^{2}-1.0+1.0)-\mu_{i}\sigma_{i}})}$
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-\mu_{i}\sigma_{i}-0.5})}.$ (18)
Notice that as $\mu_{i}-\sigma_{i}\geq 0,\sigma_{i}>0$, we must have $\mu_{i}\sigma_{i}>0$, then
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (19)
Similarly, we can show similar result for $\mu_{i}+\sigma_{i}\leq 0$. Obviously we have $\phi(\mu_{i}-\sigma_{i})\leq\phi(x^{\prime})$, and we have
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot 2\sigma_{i}\phi(\mu_{i}-\sigma_{i}))}$
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}\sigma_{i}e^{-\frac{1}{2}(\mu_{i}+\sigma_{i})^{2}})}$
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(\mu_{i}^{2}+\sigma_{i}^{2}-\log\sigma^{2}-1.0+1.0)+\mu_{i}\sigma_{i}})}$
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+\mu_{i}\sigma_{i}-0.5})}$
$\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}$ (20)
Now, consider the case when $\mu_{i}-\sigma_{i}<0<\mu_{i}+\sigma_{i}$, obviously we must have either $\phi(\mu_{i}-\sigma_{i})\leq\phi(x^{\prime})$, or $\phi(\mu_{i}+\sigma_{i})\leq\phi(x^{\prime})$. If $\phi(\mu_{i}+\sigma_{i})\leq\phi(x^{\prime})$, then the result is the same as $\mu-\sigma_{i}\geq 0$. If $\phi(\mu_{i}-\sigma_{i})\leq\phi(x^{\prime})$, then the result is the same as $\mu+\sigma_{i}\leq 0$.
Therefore, for all $\mu_{i},\sigma_{i}$, we have
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (21)
Taking the value of $K$ in, we have the result
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}$
$=\exp{(-e^{t}\cdot\sqrt{\frac{2}{\pi}}e^{-|\mu_{\max}\sigma_{\max}|-0.5})}$
$=\exp{(-e^{t}\cdot\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5})}.$ (22)
Theorem 2. The probability of quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increase exponentially to the amount of nats $t$ that codebook bitrate $\log K$ lower than the bits-back coding bitrate, i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\geq 1-e^{-t}\underbrace{\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}}_{\textup{constant}}.$ (7)
Proof.
Similar to the proof of Theorem. 1, we have
$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})=(1-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})))^{K}.$ (23)
Now we use an inequality, that $\forall y\in(0,1),K\in\mathbb{N},K\geq 1,(1-y)^{K}\geq 1-Ky$. This is due to the fact that $(1-y)^{K}$ is convex in $(0,1)$, and 1Ky1-Ky is tangent line at $y=0$. With this inequality, we have
$(1-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})))^{K}\geq 1-K(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})).$ (24)
Again, we can use integral mean value theorem, and find out that when $\mu_{i}-\sigma_{i}\geq 0$,
$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})$ $\geq 1-K(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i}))$
$\geq 1-K2\sigma_{i}\phi(\mu_{i}-\sigma_{i})$
$=1-K\sqrt{\frac{2}{\pi}}\sigma_{i}e^{-\frac{1}{2}(x_{i}-\sigma_{i})^{2}}$
$=1-K\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(x_{i}^{2}+\sigma_{i}^{2}-\log\sigma_{i}^{2}-1.0)+|\mu_{i}\sigma_{i}|-0.5}$
$=1-K\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+|\mu_{i}\sigma_{i}|-0.5}$
$\geq 1-Ke^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (25)
Similar results can be obtained for $\mu_{i}+\sigma_{i}\leq 0$. For the case that $\mu_{i}-\sigma_{i}\leq 0\leq\mu_{i}+\sigma_{i}$, we have
$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})$ $\geq 1-K(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i}))$
$\geq 1-K2\sigma_{i}\phi(0)$
$=1-K\sqrt{\frac{2}{\pi}}\sigma_{i}e^{-\frac{1}{2}(0)^{2}}$
$=1-K\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(\mu_{i}^{2}+\sigma_{i}^{2}-\log\sigma_{i}^{2}-1.0)-0.5+0.5(\mu_{i}^{2}+\sigma_{i}^{2})}$
$=1-K\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+0.5(\mu_{i}^{2}+\sigma_{i}^{2})+|\mu_{i}\sigma_{i}|-0.5}$
$=1-Ke^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (26)
Taking the value of $K=e^{D_{KL}(q(Z_{i}||\mathcal{N}(0,1)))-t}$ , we have the result
$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})$ $\geq 1-e^{D_{KL}(q(Z_{i}||\mathcal{N}(0,1)))-t-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$
$\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}.$ (27)
To better illustrate the significance of these bounds, we provide a practical example. We evaluate the ImageNet validation dataset using a pre-trained Gaussian VAE and compute that $c_{1}=8.12$ and $c_{2}=1.50$. We then visualize the upper bound and lower bound of $\textup{Pr}{|\hat{z}-\mu_{i}|\geq\sigma_{i}}$ in Fig. 3. The results show that when the codebook bitrate exceeds the bits-back coding bitrate by approximately 10 nats, the probability of large quantization error diminishes to zero. Conversely, when the codebook bitrate is smaller than the bits-back coding bitrate, the probability of large quantization error increases.
Figure 3: A visualization of large quantization error lowerbound and upperbound with ImageNet validation dataset.
Appendix B Stable Mean-KL Parametrization
We investigate an alternative to TDC, namely the Mean-KL parametrization (Lin2023MinimalRC), which is considered to be easier to train than TDC since it does not require the construction of an empirical $R(\lambda)$ model.
B.1 Mean-KL Parametrization
The Mean-KL parametrization (Lin2023MinimalRC) supports grouping with $m>1$. Its neural network output consists of two $m$-dimensional tensors, $\hat{\gamma}{i:i+m}$ and $\tau{i:i+m}$, which allocate the $D_{\mathrm{KL}}$ target $K$ across the $m$ dimensions and determine the mean, respectively. More specifically, the Mean-KL parametrization determines the mean $\mu_{i:i+m}$ and variance $\sigma^{2}_{i:i+m}$ as follows, where $\mathcal{W}(\cdot)$ denotes the principal branch of the Lambert W function:
$\gamma_{i:i+m}=\textrm{Softmax}(\hat{\gamma}_{i:i+m}),$
$\kappa_{i:i+m}=\gamma_{i:i+m}K,$
$\mu_{1:m}=\sqrt{2\kappa_{i:i+m}}\textrm{tanh}(\tau_{i:i+m}),$
$\sigma_{i:i+m}^{2}=-\mathcal{W}(-\exp(\mu_{i:i+m}^{2}-2\kappa_{i:i+m}-1.0)).$ (28)
The Mean-KL parametrization is designed for model compression. When directly applied to Gaussian VAEs, two typical cases may arise, as shown in Table 8, both of which can result in a not-a-number (NaN) error in floating-point computations.
Table 8: Two typical types of NaN in Mean-KL parametrization.
$\mu_{i}$ $\kappa_{i}$ $\sigma^{2}_{i}$
-2.7286 3.7227 NaN
[…]
B.2 Stable Mean-KL Parametrization
It is evident that the two types of NaN errors are caused by excessively large values of $|\mu_{i}|$ and excessively small values of $\kappa_{i}$, respectively. To address this numerical issue, we propose the Stable Mean-KL parametrization, which introduces two regularization parameters, $r_{1}=0.1$ and $r_{2}=0.01$. The parameter $r_{1}$ ensures that each $\kappa_{i}\geq r_{1}/m$, while $r_{2}$ shrinks $\mu_{i}$ towards 0.
$\kappa_{i:i+m}=\gamma_{i:i+m}(K-r_{1})+r_{1}/m,$
$\mu_{1:m}=\sqrt{2\kappa_{i:i+m}}\textrm{tanh}(\tau_{i:i+m})(1-r_{2}),$ (29)
B.3 Results of Stable Mean-KL Parametrization
In Table 9, we present the results of the Stable Mean-KL parametrization. For UNet-based models, Stable Mean-KL achieves performance comparable to TDC. However, for ViT-based models, Stable Mean-KL performs significantly worse than TDC. Since Stable Mean-KL does not consistently outperform TDC, we choose to use TDC for the final model. Nonetheless, if only UNet-based models are required, Stable Mean-KL can be an effective alternative to TDC, as it does not require an empirical $R(\lambda)$ model and is significantly simpler to train.
Table 9: Quantitive results on ImageNet validation dataset.
Method
bpp
(# of tokens)
UNet based ViT based
[…]
GQ (Mean-KL)
1.00
[…]
NaN NaN NaN NaN NaN NaN NaN NaN
GQ (Stable Mean-KL) 32.35 0.023 0.905 0.280 30.80 0.030 0.891 0.556
GQ (TDC) 32.47 0.023 0.907 0.322 31.71 0.024 0.903 0.349
Appendix C Implementation Details
C.1 Details of Training and Distortion Objective
We train all VQ-VAEs on the ImageNet validation dataset using $8\times$ H100 GPUs for approximately 24 hours. For UNet models, we train each model for 30 epochs using the ADAM (Kingma2014AdamAM) optimizer with a batch size of $128$ and a learning rate of $1\times 10^{-4}$. For ViT models, we train each model for 40 epochs using the ADAM optimizer with a batch size of $256$ and a learning rate of $4\times 10^{-7}$.
All VQ-VAEs are trained using the following distortion objective, which corresponds to the classical VQ-GAN (Esser2020TamingTF) objective employed in the Stable Diffusion VAE (rombach2022high).
$\Delta(X,g(z))=\mathcal{L}_{MSE}(X,g(z))+w_{1}\mathcal{L}_{LPIPS}(X,g(z))+w_{2}\mathcal{L}_{GAN}(g(z)).$ (30)
Following the implementation of Stable Diffusion, we set $w_{1}=1.0$ and $w_{2}=0.75$ for UNet models. Consistent with the implementation of BSQ (zhao2024image), we set $w_{1}=0.1$ and $w_{2}=0.1$ for ViT models.
For the image generation model, we first train all VQ-VAEs using images of size $128\times 128$, following the same settings as described above. Subsequently, we train the auto-regressive transformer for image generation using the implementation of IBQ (Shi2024TamingSV) with a Llama-base transformer architecture. The transformer has a vocabulary size of $2^{16}$, $16$ layers, $16$ attention heads, and an embedding dimension of $1024$. We train the transformer for $100$ epochs using the ADAM optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of $512$.
C.2 Details of Grouping Strategies
We extend the notation from main text. We group $m$ tokens into one large token with codebook size $K$. For each quantization output $\hat{z}_{i}$, we denote the corresponding index in the codebook as $\mathcal{I}_{i}$. And we have $c_{\mathcal{I}_{i}}=\hat{z}_{i}$.
Post Quantization (PQ) The Post Quantization (PQ) grouping strategy happens after GQ. We first train a Gaussian VAE with codebook size for each dimension $K^{1/m}$. Next, we obtain GQ tokens $\mathcal{I}_{1:d}$. Then, we group those $d$ tokens into $d/m$ groups with group size of $m$. Denote the group index as $g=0,\ldots,d/m-1$, then each group can be denoted as $\{\mathcal{I}_{gm+l}\},l=1,\ldots,m$.
In that case, we have $\max\{\mathcal{I}_{gm+l}\}\leq K^{1/m}$. Then, we can view each $\mathcal{I}_{gm+l}$ as an integer in a $K^{1/m}$-base numerical system. Then, aggregating $m$ tokens $\{\mathcal{I}_{gm+l}\}$ into one large token $I^{m}_{g}$ is as easy as concatenating $m$ tokens into a larger integer in a $K$-based numerical system: $\mathcal{I}_{g}^{m}=\sum_{l=1}^{m}\mathcal{I}_{gm+l}K^{(l-i)/m}\leq K$.
Post Training (PT) The Post Training (PT) grouping strategy happens after the training of Gaussian VAE. We still start with a Gaussian VAE with target codebook size for each dimension $K^{1/m}$. Next, instead of performing GQ for each dimension, we consider the following $m$ dimension GQ for each group $g$:
[…]
In fact, we can view one dimension GQ as a max likelihood:
$\hat{z}_{i}=\arg\max_{c_{j}\in\{c_{1:K}\}}\log q(Z_{i}=c_{j}|X),\textrm{ where }c_{1:K}\sim\mathcal{N}(0,1).$
Then we can extend the max likelihood to $m$ dimension, which is equivalent to the basic version of PT in Eq. 32.
$\hat{z}_{gm:gm+m}=\arg\max_{c_{j}\in\{c_{1:K}\}}\log q(Z_{gm:gm+m}=c_{j}|X),\textrm{ where }c_{1:K}\sim\mathcal{N}(0,I_{m}).$
When group size $m$ is large, we observe that vanilla PT in Eq. 32 leads to codebook collapse (See Table. 13) and a decay in performance. Therefore, we include a regularization term weighted by $\omega$:
[…]
Training Aware (TA) The Training Aware (TR) grouping strategy happens before the training of Gaussian VAE. More specifically, we directly interoperate grouping with the TDC, and introduce the $m$ group TDC as follows:
$\mathcal{L}_{TDC}^{m}=\sum_{g=0}^{d/m}\lambda_{g}\sum_{l=1}^{m}D_{KL}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))+\Delta(X,g(z)),$
$\textup{where }\lambda_{g}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textup{ bits},\\ \lambda_{\textrm{mean}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\in[\log_{2}K-0.5,\log_{2}K+0.5]\textup{ bits},\\ \lambda_{\textrm{max}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textup{ bits}.\end{array}\right$. (36)
Besides, the $\lambda$ update heuristic should also consider $m$ dimension as a group:
$\lambda_{\textrm{min}}=\lambda_{\textrm{min}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ min}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textrm{ else }\lambda_{\textrm{min}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{mean}}=\lambda_{\textrm{mean}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ mean}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>\log_{2}K\textrm{ else }\lambda_{\textrm{mean}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{max}}=\lambda_{\textrm{max}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ max}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textrm{ else }\lambda_{\textrm{max}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}.$ (37)
C.3 Details of Hyper-parameters
Below, we describe the implementation details along with the definition of hyperparameters for each method. In Table 10, we list the values of these hyperparameters for different bits-per-pixel (bpp) settings.
VQGAN (van2017neural) We adopt the factorized codebook VQGAN variant following (Zheng2022MoVQMQ). For each codebook, we use a codebook size of $K=2^{16}$ and a group dimension of $m=16$. The number of codebooks $n$ varies depending on the bitrate. Additionally, we use a codebook loss weight of $\lambda=1.0$ and a commitment loss weight of $\zeta=0.25$.
FSQ (mentzer2023finite) The only parameter of FSQ is the codebook list $l$, which represents the quantization level for each integer value. We set each unit value to $2^{4}=16$, and populate $l$ with the appropriate number of $16$s according to the desired bitrate.
LFQ (yu2023language) For LFQ at $0.25$ bpp, we follow the original paper and split a large codebook of size $2^{16}$ into $n=2$ smaller codebooks, each with $K=2^{8}$ and a codebook dimension of $m=8$. We use an entropy loss weight of $\lambda=0.1$ and a commitment loss weight of $\zeta=0.025$.
BSQ (zhao2024image) We fix the size of each BSQ codebook to $2^{1}$, with a group dimension of $m=1$, and vary the number of codebooks $n$ according to the desired bitrate. For the entropy penalization parameter, we set $\lambda=0.1$, following the official implementation.
GQ We use TR grouping with a fixed codebook size of $K=2^{16}$. Each group has dimension $m$, and there are $n$ groups in total, such that $m\times n=16$. The value of $m$ varies depending on the bitrate.
Table 10: Details of Hyper-parameter values.
bpp Hyper-parameters
VQ 0.25 $K=2^{16},n=1,m=16,\lambda=1.0,\zeta=0.25$
0.50 $K=2^{16},n=2,m=16,\lambda=1.0,\zeta=0.25$
1.00 $K=2^{16},n=4,m=16,\lambda=1.0,\zeta=0.25$
FSQ 0.25 $l=\{16,16,16,16\}$
0.50 $l=\{16,16,16,16,16,16,16,16\}$
1.00 $l=\{16,16,16,16,16,16,16,16,16,16,16,16,16,16,16,16\}$
LFQ 0.25 $K=2^{8},n=2,m=8,\lambda=0.1,\zeta=0.025$
0.50 $K=2^{8},n=4,m=8,\lambda=0.1,\zeta=0.025$
1.00 $K=2^{8},n=8,m=8,\lambda=0.1,\zeta=0.025$
BSQ 0.25 $K=2^{1},n=16,m=1,\lambda=0.1$
0.50 $K=2^{1},n=32,m=1,\lambda=0.1$
1.00 $K=2^{1},n=64,m=1,\lambda=0.1$
GQ 0.25 $K=2^{16},n=1,m=16,\omega=2.0$
0.50 $K=2^{16},n=2,m=8,\omega=2.0$
1.00 $K=2^{16},n=4,m=4,\omega=0.0$
Appendix D Additional Quantitive Results
D.1 Comparison to other Conversion Methods
In Table. 11, we compare our GQ to TokenBridge (wang2025bridging) and ReVQ (zhang2025quantizethenrectifyefficientvqvaetraining), which also convert a Gaussian VAE into a VQ-VAE. Those results are from their original paper. It is shown that GQ has best reconstruction metrics. Besides, only GQ has theoretical guarantee.
Table 11: Quantitive results on ImageNet dataset. Bold: best, : from paper, -: not available.
Method
Training
free
Theoretical
guarantee
bpp
(# of tokens)
[…]
OpenMagViT-V2 No No 0.07 ($2^{18}\times 256$) 21.63 0.111 0.640 1.17
TokenBridge Yes No 0.375 ($2^{6}\times 4096$) - - - 1.11
ReVQ-256T No No 0.07 ($2^{18}\times 256$) 21.96 0.121 0.640 2.05
GQ (Ours) Yes Yes 0.07 ($2^{18}\times 256$) 22.30 0.116 0.642 1.04
D.2 The TDC Parameters
In Table 12, we show the effect of different TDC parameters. It is shown that $\beta=1.001,1.01,1.1$ does not make a significant different on result. However, setting $\alpha>0.5$ is harmful to performance.
Table 12: Ablation study on TDC parameters.
[…]
0.5 1.01 27.61 0.059 0.807 0.529
0.1 1.01 27.56 0.058 0.812 0.551
1.0 1.01 27.61 0.063 0.811 0.701
0.5 1.1 27.63 0.060 0.809 0.534
0.5 1.001 27.48 0.058 0.804 0.510
D.3 Effectiveness of Grouping Strategies
In Table 13, we evaluate the effect of token grouping techniques. The scenario we consider involves grouping four $4$-bit tokens into a single $16$-bit token, which is a reasonable setting for autoregressive generation. The results show that PQ has no effect on reconstruction performance, while PT provides some improvements in PSNR and SSIM. In contrast, TR, which involves training the Gaussian VAE with a grouping target, achieves the best reconstruction performance.
Table 13: Effects of token grouping on ImageNet dataset. TR strategy has best reconstruction performance for grouping $m=4$.
[…]
Gaussian VAE (w/ TDC) no (m=1) - 4.26, 2.93-5.63 1.06 32.61 0.023 0.906 0.460
GQ no (m=1) 4 - 1.00 32.11 0.023 0.906 0.414
GQ PQ (m=4) 16 - 1.00 32.11 0.023 0.906 0.414
GQ PT (m=4) 16 - 1.00 32.15 0.023 0.907 0.428
Gaussian VAE (w/ TDC) TR (m=4) - 15.99, 14.81-17.54 1.00 32.62 0.023 0.909 0.331
GQ TR (m=4) 16 - 1.00 32.47 0.023 0.907 0.322
Additionally, in Table 14, we show the effect of the regularization parameter $\omega$ for PT and TR. For high bitrates, such as $1.00$ bpp, regularization is not required; in other words, setting $\omega=0.0$ yields the good enough codebook usage and rFID. For lower bitrates, such as $0.50$ bpp, $\omega=0.0$ leads to codebook collapse, while $\omega=2.0$ achieves the best codebook entropy and rFID.
Table 14: Ablation Study on regularization $\omega$.
[…]
0.50 $0.0$ 99.3% 14.96 30.00 0.044 0.873 0.783
$1.0$ 100.0% 15.14 30.35 0.040 0.877 0.589
$2.0$ 100.0% 15.22 30.17 0.039 0.875 0.492
$4.0$ 100.0% 14.81 28.08 0.061 0.846 1.269
1.00 $0.0$ 100.0% 15.05 32.47 0.023 0.907 0.322
$1.0$ 100.0% 15.05 32.47 0.023 0.907 0.327
$2.0$ 100.0% 15.06 32.47 0.024 0.907 0.332
$4.0$ 100.0% 15.07 32.44 0.024 0.907 0.343
Table 15: Quantitive results on COCO 2017 dataset. Bold: best.
Method
bpp
(# of tokens)
UNet based ViT based
[…]
VQGAN
0.25
[…]
26.25 0.099 0.756 14.110 25.11 0.106 0.747 11.231
FSQ 26.01 0.072 0.767 5.451 25.85 0.112 0.765 11.213
LFQ 24.60 0.164 0.722 32.789 24.46 0.143 0.729 29.975
BSQ 25.29 0.085 0.763 5.803 26.15 0.082 0.798 7.034
GQ (Ours) 27.29 0.057 0.816 3.797 27.55 0.060 0.830 5.305
VQGAN
0.50
($^{\textup{16}}\times2$048)
29.06 0.049 0.839 6.616 27.83 0.058 0.832 5.461
FSQ 29.08 0.043 0.855 4.008 28.51 0.053 0.851 5.390
LFQ 26.47 0.103 0.805 17.508 27.54 0.067 0.833 8.700
BSQ 27.58 0.057 0.844 4.465 28.19 0.049 0.858 4.587
GQ (Ours) 30.14 0.037 0.877 3.116 30.18 0.034 0.887 3.616
VQGAN
1.00
[…]
31.97 0.024 0.901 3.455 31.07 0.029 0.904 3.494
FSQ 32.30 0.022 0.917 2.797 31.48 0.023 0.911 3.045
LFQ 28.16 0.072 0.845 11.121 26.36 0.103 0.794 20.381
BSQ 30.33 0.031 0.906 2.638 31.38 0.026 0.918 2.835
GQ (Ours) 32.36 0.020 0.915 1.875 31.50 0.022 0.908 2.703
Table 16: The effect of quantization in pixel space.
[…]
$\mu_{i}=\mathbb{E}[Z_{i}|X]$ (posterior mean) 16 bits 32.92 0.020 0.913 0.46
$z_{i}\sim q(Z_{i}|X)$ (Gaussian sample) $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))=$ 4.26 bits 32.61 0.021 0.911 0.46
$\hat{z}_{i}$ (GQ) $\log_{2}K=4$ bits 32.11 0.023 0.906 0.414
D.4 The Quantization Error in Pixel Space
Previously we examine the quantization error in latent space. We can further discuss the quantization error in pixel space given the decoder is smooth. More specifically, we have:
Corollary 3. Following the setting in Theorem 1, and assuming the decoder $g(.)$ satisfy $|g(x_{1})-g(x_{2})|\leq c_{3}|x_{1}-x_{2}|$, we have:
$Pr\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq\exp(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}).$ (38)
Proof. As $|g(\hat{z}_{i})-g(\mu_{i})|\leq c_{3}|\hat{z}_{i}-\mu_{i}|$, we have $Pr\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq Pr\{c_{3}|\hat{z_{i}}-\mu_{i}|\geq c_{3}\sigma_{i}\}=Pr\{|\hat{z_{i}}-\mu_{i}|\geq\sigma_{i}\}$.
We can see that theoretically, the quantization error can be magnified by the Lipschitz constant $c_{3}$. However, we note that this is not a significant issue in practice. As shown in the Table 16, the actual loss of quality caused by GQ remains reasonable.
D.5 Robustness of Codebook to Random Seed
As the codebook is usually large enough ($2^{16}$) to compensate for the randomness. We provide additional in Table 17 showing that the random seed has little effect on reconstruction performance. We used three continuos random seeds without cherry-picking. It is clear that the performance of GQ is not affected by randomness.
Table 17: The effect of codebook randomness on the performance of GQ.
[…]
42 27.61 0.059 0.807 0.529
43 27.61 0.059 0.807 0.523
44 27.62 0.059 0.807 0.526
D.6 The Effect of Simply Increasing Codebook Size
According to Theorem 1, increasing $\log K$ too much over $D_{KL}$ will not significantly improve reconstruction but will lead to a waste of bitrate (tokens). We provide an additional in Table 18, showing that GQ trained with $14$ bits and quantized into $18$ bits does not perform as well as GQ trained with $18$ bits and quantized into $18$ bits. The drawback of simply increasing the codebook size for GQ is that it is not as effective as increasing the target of TDC to that size and quantizing with a proper $\log K=D_{KL}$.
Table 18: The effect of simply increase codebook size.
[…]
$D_{KL}(2)=14$ $\log_{2}K=14$ $2^{14}\times 1024$ 25.31 0.064 0.762 0.491
$D_{KL}(2)=14$ $\log_{2}K=18$ $2^{18}\times 1024$ 27.79 0.059 0.808 0.513
$D_{KL}(2)=18$ $\log_{2}K=18$ $2^{18}\times 1024$ 27.86 0.054 0.804 0.424
D.7 Quantized Latent Visualization
In Figure 4, we show the t-NSE (Maaten2008VisualizingDU) visualization of latent after GQ, using 5 subclass of ImageNet.
Figure 4: The t-NSE visualization of latent of GQ vs. unquantized Gaussian VAE. It is shown that the latent before and after quantization are quite similar.
D.8 More Generation Results
To better understand the generation performance, in Table 19, we present the auto-regressive generation result of GQ in different bitrate regime. And in Table 20, we present the auto-regressive generation result of GQ in FFHQ dataset. It is shown that the advantage of GQ is consistent in different bitrate regime and datasets.
Table 19: The generation performance of GQ in different bitrate regime.
Method bpp (num of tokens) gFID IS
TokenBridge $0.1875(2^{16}\times 256)$ 8.29 188.05
GQ (Ours) $0.1875(2^{16}\times 256)$ 7.74 229.53
TokenBridge $0.25(2^{16}\times 256)$ 7.82 198.24
GQ (Ours) $0.25(2^{16}\times 256)$ 7.67 230.79
Table 20: The generation performance of GQ with FFHQ dataset.
Method bpp (num of tokens) gFID IS
BSQ $0.25(2^{16}\times 256)$ 5.48 -
TokenBridge $0.25(2^{16}\times 256)$ 7.15 -
GQ (Ours) $0.25(2^{16}\times 256)$ 5.09 -
D.9 Prior-Posterior Mismatch
Sometimes the Gaussian VAE might suffer from prior-posterior mismatch. However, in our case, such mismatch is not severe. To illustrate this, we estimate the prior posterior mismatch by considering the relationship between $q(Z)$ and $\mathcal{N}(0,I)$. More specifically, we have
$D_{KL}(q(Z)||N(0,1))\approx\frac{1}{N}\sum_{i=1}^{N}(\log q(z^{i})-\log N(z^{i}|0,I)).$ (39)
Additionally, we can estimate the optimal bitrate without effected by prior posterior mismatch with a similar approximation:
$D_{KL}(q(Z|X)||p(Z))\approx\frac{1}{N}\sum_{i=1}^{N}(\log q(z^{i}|X)-\log q(z^{i})).$ (40)
We train a diffusion model to estimate $\log q(z^{i})$ by using PF-ODE and Skilling-Hutchinson trace estimator (See Appendix D.2 of Song2020ScoreBasedGM). We use the Gaussian VAE (w/ TDC) + DiT diffusion model and ImageNet validation dataset. The euler PF-ODE steps is set to 250 and the Skilling-Hutchinson number of sample is set to 1. The result is shown in Table 21. The results show that the prior-posterior mismatch is only 0.00033 bits per pixel, accounting for approximately 0.1% of the total bpp. Furthermore, the best bpp and the actual bpp show no significant difference on a scale of 0.01. This indicates that the ”bitrate waste” caused by the prior-posterior mismatch is negligible, and the mismatch itself is not significant.
Table 21: The bitrate and prior-posterior mismatch.
Divergence bits per pixel
bpp w.r.t. $\mathcal{N}(0,I)$ ($D_{KL}(q(Z\|X)\|\|\mathcal{N}(0,1))$) 0.25
bpp w.r.t. $q(Z)$ ($D_{KL}(q(Z\|X)\|\|q(Z))$) 0.25
prior posterior mismatch ($D_{KL}(q(Z)\|\|N(0,1))$) 0.000328
D.10 Complexity
Table 22: The encoding and decoding overhead of GQ over Gaussian VAE.
Method UNet based
Encoding FPS Decoding FPS
Gaussian VAE 104 64
GQ (torch) 12 61
GQ (CUDA) 79 61
As with FSQ and BSQ (mentzer2023finite; zhao2024image), our codebook can be generated on the fly by maintaining the same random number generator seed on both the encoder and decoder sides. Therefore, our GQ model has the same parameter size as the vanilla Gaussian VAE. In Table 22, we compare the encoding and decoding frames per second (FPS) of the Gaussian VAE and GQ. We use $256\times 256$ images with a batch size of $1$, and we report the wall clock time, meaning that the time required for loading data is included. The results show that the encoding FPS of GQ (implemented in PyTorch) is 12 on an H100 GPU, which is considerably slower than the 104 FPS achieved by the Gaussian VAE. On the other hand, GQ does not incur any decoding overhead.
To reduce the computational complexity of GQ, we implement GQ using a tailored CUDA kernel. Specifically, we follow the approach of Vonderfecht2025LossyCW, with a key difference: we maintain the codebook, as our bottleneck is not codebook instantiation. Additionally, we avoid the creation of large buffer vectors by performing the summation over $m$ within the CUDA kernel instead of in PyTorch. With this approach, we achieve an encoding FPS of approximately 80, with negligible overhead compared to the Gaussian VAE. A detailed comparison between the PyTorch implementation and the CUDA implementation of GQ is provided below as GQ_torch and GQ_CUDA, respectively.
1def GQ_torch(mu, sigma, codebook, m, bs, K):
2 # mu.shape = (bs, m)
3 # sigma.shape = (bs, m)
4 # codebook.shape = (K, m)
5
6 # This step create (bs, m, K) vector, which is the performance bottleneck
7 dist_m =((mu[:,None] - codebook[None])/sigma[:,None])**2
8 dist = torch.sum(dist_m, dim=1) # sum over m dimension
9 indices = torch.argmin(dist, dim=1) # argmax over K dimension
10 zhat = torch.index_select(codebook, 0, indices) # select quantized results
11 return indices, zhat
1def GQ_CUDA(mu, sigma, codebook, m, bs, K):
2 dist = torch.zeros([bs, K])
3 # need an extension wrapping and register the kernel into operator, we omit it in paper
4 # see code appendix for details
5 GQ_Kernel<<<bs * K / 256,256>>>(mu, sigma, codebook, dist, m, bs, K)
6
7 indices = torch.argmin(dist, dim=1) # argmax over K dimension
8 zhat = torch.index_select(codebook, 0, indices) # select quantized results
9 return indices, zhat
10
11__global__ void GQ_Kernel(
12 const float* mu,
13 const float* sigma,
14 const float* codebook,
15 float* dist,
16 int64_t m,
17 int64_t bs,
18 int64_t K
19) {
20 int idx = blockIdx.x * blockDim.x + threadIdx.x;
21 if (idx >= K * bs) return;
22 int bi = idx / K;
23 int ni = idx % K;
24 float a = 0.0f;
25 for (int i = 0; i < m; i++) {
26 float b = (codebook[ni * m + i] - mu[bi * dim + i]) / sigma[bi * dim + i];
27 a += b * b;
28 }
29 dist[idx] = a;
30 return;
31}
D.11 Asymptotic Complexity
It is noteworthy that GQ without grouping quantization, such as PT or TR, or GQ with a group size of $m=1$, is asymptotically faster than reverse channel coding methods. This is because, for $m=1$, the GQ target in Eq.3 reduces to a quadratic form. In this case, it suffices to sort the scalar codebook $c_{1:K}$ in advance. Despite the sorting takes $\Omega(D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)))$, the sorting is only need to be done once and can be amortized across dimension and dataset. Subsequently, the minimization in Eq.3 can be performed in $O(D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ time using binary search. The details is shown in Algorithm 2.
On the other hand, most reverse channel coding methods require $O(2^{D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1))})$ computational complexity (havasi2018minimal; Flamich2020CompressingIB; Theis2021AlgorithmsFT). A coding (Flamich2022FastRE) can achieve $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ encoding complexity, albeit at the cost of increased decoding complexity.
However, we note that this complexity advantage is not particularly meaningful in practice. This is because any auto-regressive generation model requires a softmax operation over the entire codebook, which has a complexity of $O(2^{D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1))})$. In practice, only tractable codebook sizes, such as $2^{16}$ or $2^{18}$, are used.
Appendix E Additional Quantitative Results
E.1 Additional Qualitative Results and Failure Cases
In Figure 5, we present additional qualitative results showing that GQ achieves superior visual quality. However, we also note that none of the approaches is successful in reconstructing the plate of the residential vehicle. The text content remains challenging for low bitrate VQ-VAEs.
1input $c_{1:K}(\textrm{sorted, such that }c_{j}\leq c_{j+1}),\mu_{i}$
[…]
3for $j=1$ to $K$ do
[…]
6return $c_{j^{*}},j^{*}$
Algorithm 1 GQ (argmax)
1input $c_{1:K}(\textrm{sorted, such that }c_{j}\leq c_{j+1}),\mu_{i}$
2l=1,r=Kl=1,r=K
3while $l+1<r$ do
4   $m=(l+r)//2$
5   if $c_{m}<\mu_{i}$ do
6    $l=m$
7   else
8    $r=m$
[…]
10   return $c_{l},l$
11else
12   return $c_{r},r$
Algorithm 2 GQ (bisect)
Figure 5: Qualitative results on ImageNet dataset and 0.25 bpp. None of those approaches correctly reconstruct the plate number.

### v2

@@
 Abstract
-Vector quantized variational autoencoder (VQ-VAE) is a discrete auto-encoder that compresses images into discrete tokens.
-It is difficult to train due to discretization.
-In this paper, we propose a simple yet effective technique, dubbed Gaussian Quant (GQ), that converts a Gaussian VAE with certain constraint into a VQ-VAE without training.
-GQ generates random Gaussian noise as a codebook and finds the closest noise to the posterior mean.
+Vector-quantized variational autoencoders (VQ-VAEs) are discrete autoencoders that compress images into discrete tokens.
+However, they are difficult to train due to discretization.
+In this paper, we propose a simple yet effective technique dubbed Gaussian Quant (GQ), which first trains a Gaussian VAE under certain constraints and then converts it into a VQ-VAE without additional training.
+For conversion, GQ generates random Gaussian noise as a codebook and finds the closest noise vector to the posterior mean.
 Theoretically, we prove that when the logarithm of the codebook size exceeds the bits-back coding rate of the Gaussian VAE, a small quantization error is guaranteed.
-Practically, we propose a heuristic to train Gaussian VAE for effective GQ, named target divergence constraint (TDC).
+Practically, we propose a heuristic to train Gaussian VAEs for effective conversion, named the target divergence constraint (TDC).
 Empirically, we show that GQ outperforms previous VQ-VAEs, such as VQGAN, FSQ, LFQ, and BSQ, on both UNet and ViT architectures.
-Furthermore, TDC also improves upon previous Gaussian VAE discretization methods, such as TokenBridge.
-The source code is provided in https://github.com/tongdaxu/VQ-VAE-from-Gaussian-VAE.
+Furthermore, TDC also improves previous Gaussian VAE discretization methods, such as TokenBridge.
+The source code is provided in the supplementary materials.
+Machine Learning, ICML
+Figure 1: The rate-distortion performance on the ImageNet dataset demonstrates that GQ outperforms previous VQ-VAEs on both UNet and ViT architectures.
 1 Introduction
-Vector-quantized variational autoencoder (van2017neural) is an autoencoder that compresses images into discrete tokens.
-It is fundamental to autoregressive generative models (esser2021taming; chang2022maskgit; yu2023language; sun2024autoregressive).
-However, VQ-VAE is difficult to train: the encoding process of VQ-VAE is not differentiable and challenges such as codebook collapse often emerge (sonderby2017continuous).
-Special techniques are required to ensure the convergence of VQ-VAE, such as commitment loss (van2017neural), expectation maximization (EM) (roy2018theory), Gumbel-Softmax (jang2016categorical; maddison2016concrete; sonderby2017continuous), and entropy loss (yu2023language; zhao2024image).
-In this paper, we circumvent the challenge of training VQ-VAE by converting a Gaussian VAE with certain constraint into a VQ-VAE without any training.
-More specifically, we propose Gaussian Quant (GQ), a simple yet effective method for training-free conversion.
-The core idea is to generate a codebook of one-dimensional Gaussian noise and, for each dimension of the posterior, select the codebook entry that is closest to the posterior mean.
-Theoretically, we show that as the logarithm of the codebook size exceeds the bits-back coding bitrate (hinton1993keeping; townsend2019practical) of the Gaussian VAE, the resulting quantization error is small.
-In other words, GQ and the Gaussian VAE exhibit similar rate-distortion performance.
-This result serves as the theoretical foundation of GQ and provides a principled guideline for selecting codebook sizes.
-Practically, we introduce the target divergence constraint (TDC) to train a Gaussian VAE for efficient conversion.
-TDC encourages the Gaussian VAE to achieve the same Kullback–Leibler (KL) divergence for each dimension, corresponding to the bits-back coding bitrate.
-Empirically, we demonstrate that GQ with Gaussian VAE trained by TDC, outperforms previous VQ-VAEs such as VQGAN, FSQ, LFQ, and BSQ (van2017neural; mentzer2023finite; yu2023language; zhao2024image) in terms of reconstruction quality, using both UNet and ViT backbones.
-Additionally, we show that TDC can improve previous Gaussian VAE discretization methods, such as TokenBridge (wang2025bridging).
-Figure 1: The rate-distortion performance on the ImageNet dataset demonstrates that GQ outperforms previous VQ-VAEs on both UNet and ViT architectures.
-Our contributions can be summarized as follows:
-  • (Section 3.1) We propose GQ, a simple yet effective approach that converts a pre-trained Gaussian VAE with certain constraint into VQ-VAE without training.
-  • (Section 3.2) Theoretically, we prove that when the codebook size of GQ is close to the bits-back coding bitrate of the Gaussian VAE, the conversion error remains small.
-  • (Section 3.3) Empirically, we introduce target divergence constraint (TDC) to implement GQ and show that GQ outperforms previous VQ-VAEs, such as VQGAN, FSQ, LFQ, and BSQ, on both UNet and ViT architectures.
-  • (Section 3.5) Furthermore, we show that TDC can be used to improve previous Gaussian VAE discretization approaches, such as TokenBridge.
+Vector-quantized variational autoencoders (VQ-VAEs) (Van Den Oord et al., 2017) compress images into discrete tokens and are fundamental to autoregressive generative models (Esser et al., 2021; Chang et al., 2022; Yu et al., 2023; Sun et al., 2024a).
+However, VQ-VAEs are difficult to train: the encoding process is non-differentiable, and issues such as codebook collapse often arise (Sønderby et al., 2017).
+As a result, special techniques are required, including commitment loss (Van Den Oord et al., 2017), expectation maximization (EM) (Roy et al., 2018), Gumbel-Softmax (Jang et al., 2016; Maddison et al., 2016; Sønderby et al., 2017), and entropy loss (Yu et al., 2023; Zhao et al., 2024).
+Figure 2: Qualitative results on ImageNet dataset at 0.25 bits-per-pixel (bpp). GQ has the most visually pleasing reconstruction.
+In this paper, we circumvent the difficulty of training VQ-VAEs by not training them at all.
+Specifically, we propose Gaussian Quant (GQ), which first trains a constrained Gaussian VAE and then converts it into a VQ-VAE without additional training.
+For conversion, GQ generates a codebook of one-dimensional Gaussian noise and, for each posterior dimension, select the entry closest to the posterior mean.
+We show theoretically that when the logarithm of the codebook size exceeds the bits-back coding rate (Hinton and Van Camp, 1993; Townsend et al., 2019) of the Gaussian VAE, the quantization error is small, providing a principled guideline for codebook size selection.
+To train a Gaussian VAE for effective conversion, we introduce the target divergence constraint (TDC), which encourages the Gaussian VAE to achieve the same Kullback–Leibler (KL) divergence for each dimension, corresponding to the bits-back coding bitrate.
+Empirically, we demonstrate that GQ with Gaussian VAE trained by TDC, outperforms previous VQ-VAEs such as VQGAN, FSQ, LFQ, and BSQ (Van Den Oord et al., 2017; Mentzer et al., 2023; Yu et al., 2023; Zhao et al., 2024) in terms of reconstruction, on UNet and ViT backbones.
+Additionally, we show that TDC improves previous Gaussian VAE discretization methods, such as TokenBridge (Wang et al., 2025).
+Our contributions are summarized as follows:
+  • (Section 3.1) We propose Gaussian Quant (GQ), a simple yet effective approach that constructs a VQ-VAE by training a constrained Gaussian VAE and converting it into a VQ-VAE without additional training.
+  • (Section 3.2) We theoretically show that when the GQ codebook size matches the bits-back coding rate of the Gaussian VAE, the conversion error remains small.
+  • (Section 3.3) We introduce the target divergence constraint (TDC) to train Gaussian VAEs for effective conversion, and show that GQ outperforms prior VQ-VAEs on both UNet and ViT architectures.
+  • (Section 3.6) We show that TDC improves TokenBridge, an existing Gaussian VAE conversion methods.
 2 Preliminaries
-We denote the original image as $X$, the latent variable as $Z$, the encoder as $f(\cdot)$, and the decoder as $g(\cdot)$.
-We use $\log(\cdot)$ to denote the natural logarithm (base $e$) and KL divergence as $D_{KL}(\cdot||\cdot)$.
-Similarly, we use $\log_{2}(\cdot)$ to denote the logarithm with base $2$, and KL divergence as $D_{KL(2)}(\cdot||\cdot)$.
+We denote the original image as $X$, the latent as $Z$, the encoder as $f(\cdot)$, and the decoder as $g(\cdot)$.
+We use $\log(\cdot)$ for the natural logarithm and $D_{KL}(\cdot||\cdot)$ for KL divergence.
+We denote the bits-back coding bitrate as $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,I))$.
+We denote the base-2 logarithm as $\log_{2}(\cdot)$, with the corresponding bitrate $R_{i}^{\mathrm{bits}}=R_{i}/\log 2$.
 2.1 Vector Quantized Variational Autoencoder
-VQ-VAE (van2017neural) transforms source image into a series of integer tokens, which can be decoded using a codebook and decoder.
-To facilitate auto-regressive generation, it typically involves a deterministic transformation and a shared codebook across different tokens.
-More specifically, VQ-VAE maintains a codebook $c_{1:K}$ with size $K$ and a bitrate of $\log K$.
-The encoding process of VQ-VAE involves finding the closest codeword $c_{j}$ in $c_{1:K}$ to the encoder output $f(x)_{i}$ for each latent dimension $i$.
-Denote distortion as $\Delta(\cdot,\cdot)$, the optimization target of VQ-VAE is the rate-distortion function weighted by Lagrangian multiplier $\lambda$:
-$\mathcal{L}_{VQ}=\lambda\underbrace{\log K}_{\textrm{bitrate}}+\underbrace{\mathbb{E}[\Delta(X,g(\hat{z}))]}_{\textrm{distortion}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{Reg}},$
+VQ-VAE (Van Den Oord et al., 2017) encodes images into integer tokens.
+For autoregressive generation, it uses deterministic encoding and a shared codebook.
+Specifically, VQ-VAE maintains a codebook $c_{1:K}$ of size $K$ with bitrate $\log K$.
+For each dimension $i$, the encoder output $f(x)_{i}$ is quantized to the closest codeword $c_{j}$ in $c_{1:K}$.
+Denoting distortion as $\Delta(\cdot,\cdot)$, the VQ-VAE objective is the rate-distortion function weighted by a Lagrange multiplier $\lambda$:
+$\mathcal{L}_{VQ}=\lambda\underbrace{\log K}_{\textrm{bitrate}}+\underbrace{\mathbb{E}[\Delta(X,g(\hat{z}))]}_{\textrm{distortion}}+\mathcal{L}_{Reg},$
 […]
-and $\mathcal{L}_{Reg}$ are regularization for VQ-VAE to converge, such as a combination of commitment loss and codebook loss (van2017neural) and Gumbel Softmax loss (sonderby2017continuous).
-2.2 Gaussian Variational Autoencoder and Bits-Back Coding
-The Gaussian VAE is a special type of VAE (kingma2013auto) with a prior $\mathcal{N}(0,I)$ and a fully factorized Gaussian posterior $q(Z|X)$.
-The encoding process of a Gaussian VAE simply involves sampling the latent variable $z_{i}\sim q(Z_{i}|X)$ for each latent dimension $i$.
-Assuming $\log p(X|Z=z)\propto(1/\lambda)\Delta(X,g(z))$, then the negative evidence lower bound (ELBO) of the Gaussian VAE is equivalent to a rate-distortion function of a bits-back coding bitrate term and a distortion term:
+where $c_{1:K}$ is the learned codebook, and $\mathcal{L}_{\mathrm{Reg}}$ is a regularization term that ensures VQ-VAE convergence, such as commitment loss, codebook loss (Van Den Oord et al., 2017), and Gumbel-Softmax loss (Sønderby et al., 2017).
+2.2 Gaussian VAE and Bits-Back Coding
+The Gaussian VAE is a special VAE (Kingma et al., 2013) with a prior $\mathcal{N}(0,I)$ and a fully factorized Gaussian posterior $q(Z|X)$.
+Its encoding simply involves sampling each latent dimension $z_{i}\sim q(Z_{i}|X)$.
+Assuming $\log p(X|Z=z)\propto(1/\lambda)\Delta(X,g(z))$, the negative evidence lower bound (ELBO) of the Gaussian VAE corresponds to a rate-distortion function, consisting of a bits-back coding bitrate term and a distortion term:
 $\mathcal{L}_{VAE}=\lambda\underbrace{D_{KL}(q(Z|X)||\mathcal{N}(0,1))}_{\textrm{bits-back coding bitrate}}+\underbrace{\mathbb{E}[\Delta(X,g(z))]}_{\textrm{distortion}},$
 $z_{i}\sim q(Z_{i}|X=x)=\mathcal{N}(\mu_{i},\sigma_{i}^{2}),i=1\ldots d.$ (2)
-The bitrate of $z_{i}$ is the bits-back coding bitrate, defined as $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,I))$ (hinton1993keeping; townsend2019practical).
-This is because, when compressing $X$ losslessly, one can communicate $z_{i}$ using $D_{KL}(q(Z|X)||\mathcal{N}(0,I))$ nats for arbitrary precision.
+The bits-back coding bitrate of $Z_{i}$ is defined as $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,I))$ (Hinton and Van Camp, 1993), because when compressing $X$ losslessly, $z_{i}$ can be communicated using $R_{i}$ nats to arbitrary precision.
 3 Gaussian Quant: Vector Quantization using Gaussian VAE
-3.1 Direct Quantization of Gaussian VAE
-We propose an extremely simple technique to obtain a VQ-VAE from a Gaussian VAE: we directly generate one-dimensional Gaussian noise as the codebook for VQ-VAE (van2017neural) and quantize the posterior mean $\mu_{i}$ of the Gaussian VAE independently for each dimension $i$.
-Because the codebook consists entirely of samples from a Gaussian distribution, we refer to our approach as Gaussian Quant (GQ).
-Specifically, we randomly generate $K$ codebook values $c_{1:K}\sim\mathcal{N}(0,1)$, which is the same for each dimension.
-Then, for each dimension $i$, we select the $c_{j}$ that is closest to the posterior mean $\mu_{i}$ and denote the quantized value as $\hat{z}_{i}$:
+We propose an alternative approach to obtain a VQ-VAE: first train a constrained Gaussian VAE, then convert it into a VQ-VAE.
+3.1 Converting a Gaussian VAE into a VQ-VAE
+Given a Gaussian VAE, we convert it into a VQ-VAE by generating one-dimensional Gaussian noise as a codebook and quantizing the posterior mean independently for each latent dimension.
+The codebook is fixed once generated.
+Because it consists entirely of Gaussian samples, we call this approach Gaussian Quant (GQ).
+Specifically, we randomly generate a codebook $c_{1:K}\sim\mathcal{N}(0,1)$ shared across all dimensions.
+Then, for each dimension $i$, we select the codeword $c_{j}$ closest to the posterior mean $\mu_{i}$ and denote the quantized value as $\hat{z}_{i}$:
 […]
 3.2 Theoretical Relationship between the Codebook Size and Quantization Error
-Why GQ works and how to select $K$ are not straightforward questions.
-Theoretically, we show that GQ preserves the rate-distortion property of the Gaussian VAE: when the bitrate $\log K$ matches the bits-back coding bitrate of the Gaussian VAE, the quantization error is small.
-More specifically, we show that the probability of large quantization error decays doubly exponentially as the codebook bitrate $\log K$ exceeds the bits-back coding bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$.
-Theorem 1.
+Why GQ works and how to choose $K$ are not obvious.
+Theoretically, we show that GQ preserves the rate-distortion properties of the Gaussian VAE: when the codebook bitrate $\log K$ matches the bits-back coding rate $R_{i}$ of the Gaussian VAE, the quantization error is small.
+More specifically, we denote the event $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ as large quantization error.
+Then, we show that the probability of a large quantization error decays doubly exponentially as the codebook bitrate $\log K$ exceeds the bits-back coding rate $R_{i}$.
+Theorem 3.1.
 Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i}$ and $\sigma_{i}$, respectively.
-Assuming that the product and sum satisfy $|\mu_{i}\sigma_{i}|\leq c_{1}$ and $|\mu_{i}|+|\sigma_{i}|\leq c_{2}$, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays doubly exponentially with respect to the number of nats $t$ by which the codebook bitrate $\log K$ exceeds the bits-back coding bitrate.
-i.e.,
-$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+t,$
-$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-e^{t}\underbrace{\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}}_{\textup{constant}})}.$ (4)
-Conversely, when the codebook bitrate $\log K$ is smaller than the bits-back coding bitrate, the probability of large quantization error increases exponentially toward $1$.
-More specifically, we show that the probability of large quantization error increases exponentially when the codebook bitrate $\log K$ is lower than the bits-back bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$.
-Theorem 2.
-The probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increases exponentially with respect to the number of nats $t$ by which the codebook bitrate $\log K$ is lower than the bits-back coding bitrate, i.e.,
-$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-t,$
-$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\geq 1-e^{-t}\underbrace{\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}}_{\textup{constant}}.$ (5)
-Theorems 1 and 2 provide a principled guideline for choosing $K$, such that $\log K$ should be close to the bits-back bitrate $D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$.
-In practice, setting $\log_{2}K=\left\lceil D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\right\rfloor$ typically yields a small enough reconstruction error, where $\left\lceil\cdot\right\rfloor$ denotes the rounding operator.
-Using a larger $K$ does not provide additional benefits, while a smaller $K$ increases the error significantly.
-3.3 Practical Implementation with Target Divergence Constraint
-There are two challenges to make GQ practical.
-The first challenge is, if we want to construct a VQ-VAE with a specific codebook size $K$, how can we train a Gaussian VAE with corresponding KL divergence?
-The second challenge is, for a vanilla Gaussian VAE trained to minimize the loss in Eq.2, the values of $D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))$ vary significantly across dimensions $i$.
-How to train a Gaussian VAE with KL divergence remains close to $\log K$ across each dimension?
+We assume $|\mu_{i}\sigma_{i}|\leq c_{1}$ and $|\mu_{i}|+|\sigma_{i}|\leq c_{2}$.
+Given a fixed $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays doubly exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ exceeds the bits-back coding rate.
+That is,
+$\textup{when }\log K=R_{i}+t,$
+$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\leq\exp{(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5})}.$ (4)
+Conversely, when the codebook bitrate $\log K$ is smaller than the bits-back coding rate $R_{i}$, the probability of a large quantization error increases exponentially toward $1$.
+Theorem 3.2.
+Following Theorem 3.1, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increases exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ falls below the bits-back coding rate.
+That is,
+$\textup{when }\log K=R_{i}-t,$
+$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}.$ (5)
+Theorems 3.1 and 3.2 provide a principled guideline for choosing $K$: $\log K$ should be close to the bits-back rate $R_{i}$.
+In practice, setting $\log_{2}K=\left\lceil R_{i}^{\mathrm{bits}}\right\rfloor$ typically yields sufficiently small reconstruction error, where $\left\lceil\cdot\right\rfloor$ denotes rounding.
+Using a larger $K$ offers no additional benefit, while a smaller $K$ significantly increases the error.
+3.3 Training a Conversion-ready Gaussian VAE with Target Divergence Constraint
+Now that we know how to convert a Gaussian VAE into a VQ-VAE, we discuss how to train a Gaussian VAE for effective conversion.
+To construct a VQ-VAE with a specific codebook size $K$, we need to train the Gaussian VAE so that $R_{i}$ is close to $\log K$ for all dimensions $i=1,\dots,d$.
+To train a Gaussian VAE under this constraint, we propose a heuristic Target Divergence Constraint (TDC).
+Specifically, we set the target $R_{i}$ to $\log K$.
+For each dimension $i$, we apply a larger penalty if $R_{i}$ exceeds $\log K+\alpha$ and a smaller penalty if $R_{i}$ falls below $\log K-\alpha$, using different $\lambda$ values for each case.
+Here, $\alpha$ is a hyperparameter controlling the threshold:
+$\mathcal{L}_{TDC}=\sum_{i=1}^{d}\lambda_{i}R_{i}+\mathbb{E}[\Delta(X,g(z))],$
+$\textup{where }\lambda_{i}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&R_{i}<\log K-\alpha,\\ \lambda_{\textrm{mean}},&R_{i}\in[\log K\pm\alpha],\\ \lambda_{\textrm{max}},&R_{i}>\log K+\alpha.\end{array}\right$.
+(9)
+To determine $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$, we initialize them to 1 and update them according to the relationship between the statistics of $R_{i}$ and $\log K\pm\alpha$ during each gradient descent step:
+$\lambda_{\textrm{min}}=\beta\lambda_{\textrm{min}}\textrm{ if}\textrm{ min}_{i}\{R_{i}\}>\log K-\alpha\textrm{ else }\beta^{-1}\lambda_{\textrm{min}},$
+$\lambda_{\textrm{mean}}=\beta\lambda_{\textrm{mean}}\textrm{ if}\textrm{ mean}_{i}\{R_{i}\}>\log K\textrm{ else }\beta^{-1}\lambda_{\textrm{mean}},$
+$\lambda_{\textrm{max}}=\beta\lambda_{\textrm{max}}\textrm{ if}\textrm{ max}_{i}\{R_{i}\}>\log K+\alpha\textrm{ else }\beta^{-1}\lambda_{\textrm{max}},$
+where $\beta$ is a hyperparameter controlling the update speed.
+To avoid numerical issues, we clip $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$ to the range $10^{-3},10^{3}]$ after each update.
+3.4 Supporting Multi-dimensional Codebook
+The vanilla VQ-VAE has three key parameters: codebook size $K$, codebook dimension $m$, and number of tokens $N$.
+Now GQ is able to construct VQ-VAE with codebook dimension $m=1$, we can extend GQ to codebook dimension $m>1$.
+Specifically, given a Gaussian VAE, we can quantize $m$ latents into a single large token with codebook size $\log_{2}K=\left\lceil\sum_{k=i}^{i+m}R_{k}^{\mathrm{bits}}\right\rfloor$.
+During quantization, we minimize the $\sigma_{i}$-weighted distance to find nearest neighbour, with the codebook $c_{1:K}$ sampled from $m$ dimensional factorized standard Gaussian:
 […]
-$\mathcal{L}_{TDC}=\sum_{i=1}^{d}\lambda_{i}D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+\Delta(X,g(z)),$
-$\textup{where }\lambda_{i}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textup{ bits},\\ \lambda_{\textrm{mean}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\in[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha,\log_{2}K+\alpha}]\textup{ bits},\\ \lambda_{\textrm{max}},&D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textup{ bits}.\end{array}\right$.
+ $\textrm{ where }c_{1:K}\sim\mathcal{N}(0,I_{m}).$ (10)
+However, for low-bitrate cases, this vanilla $m$-dimensional conversion leads to codebook collapse.
+This occurs because $|\mu_{i}|$ is bounded by $\sqrt{2R_{i}}$ (see Lemma A.1), so when $R_{i}$ is small, codebook vectors far from 0 are never selected.
+To address this, we introduce a regularization term weighted by $\omega$ that encourages the selection of $c_{j}$ farther from 0:
+[…]
+Now we have the conversion method for multi-dimensional codebook, we consider how to train a Gaussian VAE for multi-dimensional codebook.
+Specifically, we modify the TDC training target to account for the relationship between the sum of $m$ bitrates $\sum_{k=i}^{i+m}R_{k}$ and $\log_{2}K\pm\alpha$.
+The detailed TDC update rule is:
+$\mathcal{L}_{TDC}^{m}=\sum_{j=1}^{d//m}\lambda_{j}\sum_{k=jm}^{jm+m}R_{k}+\mathbb{E}[\Delta(X,g(z))],$
+$\textup{where }\lambda_{i}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&\sum_{k=jm}^{jm+m}R_{k}<\log_{2}K-\alpha,\\ \lambda_{\textrm{mean}},&\sum_{k=jm}^{jm+m}R_{k}\in[\log_{2}K\pm\alpha],\\ \lambda_{\textrm{max}},&\sum_{k=jm}^{jm+m}R_{k}>\log_{2}K+\alpha.\end{array}\right$.
 (14)
-To determine $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$, we extend the heuristic in MIRACLE (havasi2018minimal) and HiFiC (Mentzer2020HighFidelityGI).
-More specifically, we initialize $\lambda_{\textrm{min}}=\lambda_{\textrm{mean}}=\lambda_{\textrm{max}}=1$, and update them according to the following rule:
-$\lambda_{\textrm{min}}=\lambda_{\textrm{min}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ min}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textrm{ else }\lambda_{\textrm{min}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
-$\lambda_{\textrm{mean}}=\lambda_{\textrm{mean}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ mean}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>\log_{2}K\textrm{ else }\lambda_{\textrm{mean}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
-$\lambda_{\textrm{max}}=\lambda_{\textrm{max}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ max}_{i}\{D_{KL(2)}(q(Z_{i}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textrm{ else }\lambda_{\textrm{max}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$ (10)
-where $\beta$ is the hyper-parameter controlling update speed.
-To avoid numerical issue, we further clip $\lambda_{\textrm{min}},\lambda_{\textrm{mean}},\lambda_{\textrm{max}}$ into a range of $10^{-3},10^{3}]$ after each update.
-In practice, we use $\alpha=0.5,\beta=1.01$.
-In Appendix B, we propose an alternative implementation of TDC using Lambert W function, which is less effective for ViT models.
-3.4 Grouping to Multiple Dimensions
-The vanilla VQ-VAE has three key parameters: codebook size, codebook dimension and number of token.
-Some previous alternative to VQ-VAE, such as LFQ and BSQ (yu2023language; zhao2024image) limit the codebook dimension to 1.
-To support codebook dimension greater than 11, we can group $m$ tokens into a single large token with codebook size $\log_{2}K=\left\lceil\sum_{l=i}^{i+m}D_{KL}(q(Z_{l}|X)||\mathcal{N}(0,1))\right\rfloor$.
-There are three grouping strategies to achieve this: post-quantization (PQ), post-training (PT), and training-based (TR), with different trade-offs in flexibility and performance.
-We brief these strategies in main text and provide details in Appendix C.2.
-PQ is the most flexible grouping strategy and can be applied after GQ.
-For PQ, since the posterior of the Gaussian VAE $q(Z|X)$ is a factorized Gaussian, the quantization is also independent across dimensions.
-This means that we can trivially combine $m$ tokens into a larger one, by treating each token as an integer in $K^{1/m}$-based number system and aggregate them, which is the same as other one dimensional VQ-VAEs (chang2022maskgit; mentzer2023finite; zhao2024image).
-PT is less flexible than PQ, as it can only be applied before GQ and after training the Gaussian VAE.
-For PT, we can view the one-dimensional GQ in Eq.
-3 as the maximum likelihood estimator of a one-dimensional Gaussian.
-This approach can be extended to an $m$-dimensional diagonal Gaussian distribution.
-Additionally, for low bitrate cases, we observe that $m$-dimensional PT leads to low codebook usage.
-This is because $|\mu_{i}|$ is bounded by $\sqrt{2D_{KL}}$.
-When the $D_{KL}$ is small, some vector in the codebook that is far from 0 is never used.
-To address this, we introduce a regularization term weighted by $\omega$ to improve codebook usage by encouraging the selection $c_{j}$ that is far from 0:
 […]
-TR is the least flexible strategy and must be used during the training of the Gaussian VAE.
-Specifically, building on the PT quantization approach, we can relax the TDC training target by considering the relationship between $\sum_{l=i}^{i+m}D_{KL}(q(Z_{l}|X)||\mathcal{N}(0,1))$ and $\log_{2}K\pm 0.5$ bits.
-In terms of performance, PQ does not affect reconstruction at all.
-PT provides slight improvements in reconstruction, while TR significantly enhances reconstruction performance (see Table 13).
+Table 1: Quantitative results on the ImageNet dataset, comparing different VQ methods using the same model architecture.
+Our GQ outperforms other VQ methods on both UNet and ViT architectures.
+Bold: best.
+Gray: continuous Gaussian VAE.
+Method bpp ($K\times N$) UNet based ViT based
+[…]
+Gaussian VAE $\approx$ 0.25 (-) 28.60 0.047 0.849 0.556 28.95 0.045 0.851 0.672
+Gaussian VAE (w/ TDC) 28.05 0.053 0.829 0.535 28.44 0.054 0.837 0.793
+[…]
+FSQ (Mentzer et al., 2023) 26.34 0.075 0.756 1.125 25.87 0.109 0.751 3.856
+LFQ (Yu et al., 2023) 24.74 0.164 0.722 16.337 24.81 0.143 0.725 15.71
+BSQ (Zhao et al., 2024) 25.62 0.086 0.754 1.080 26.52 0.083 0.793 1.649
+GQ (Ours) 27.61 0.059 0.807 0.529 27.88 0.061 0.823 0.932
+Gaussian VAE $\approx$ 0.50 (-) 31.18 0.028 0.889 0.389 31.64 0.030 0.900 0.867
+Gaussian VAE (w/ TDC) 30.84 0.035 0.884 0.549 31.18 0.033 0.896 0.757
+VQGAN (Esser et al., 2021) 0.50 ($^{\textup{16}}\times2$048) 29.21 0.052 0.831 1.600 27.86 0.062 0.823 1.228
+FSQ (Mentzer et al., 2023) 29.29 0.047 0.845 0.871 28.83 0.055 0.842 1.067
+LFQ (Yu et al., 2023) 26.90 0.107 0.800 8.035 27.87 0.068 0.829 2.444
+BSQ (Zhao et al., 2024) 27.88 0.059 0.836 0.788 28.44 0.051 0.852 0.700
+GQ (Ours) 30.17 0.039 0.875 0.492 30.42 0.037 0.882 0.592
+Gaussian VAE $\approx$ 1.00 (-) 32.73 0.022 0.910 0.490 32.28 0.023 0.910 0.469
+Gaussian VAE (w/ TDC) 32.75 0.021 0.910 0.294 32.05 0.023 0.909 0.397
+[…]
+FSQ (Mentzer et al., 2023) 32.38 0.025 0.905 0.636 31.58 0.026 0.905 0.544
+LFQ (Yu et al., 2023) 28.31 0.074 0.840 3.617 26.67 0.105 0.790 8.288
+BSQ (Zhao et al., 2024) 30.50 0.032 0.900 0.346 31.60 0.027 0.914 0.379
+GQ (Ours) 32.47 0.023 0.907 0.322 31.71 0.024 0.903 0.349
+3.5 Practical Guidelines
+To summarize, GQ has the same interface as a VQ-VAE and constructs a VQ-VAE with codebook size $K$, codebook dimension $m$, and number of tokens $N$ with two steps:
+  • First, GQ trains a Gaussian VAE with latent dimension $d=mN$, using the TDC constraint with target $\log K$ as in Eq. 14.
+  • Next, GQ converts this constrained Gaussian VAE into a VQ-VAE using Eq. 10. The codebook $c_{1:K}$ is fixed once generated.
 3.6 Improving TokenBridge with Target Divergence Constraint
-TokenBridge (wang2025bridging) also convert a pre-trained Gaussian VAE into a VQ-VAE.
-It adopts the Post Training Quantization (PTQ) technique from model compression and proposes to treat latent as model parameters to discretize.
-It uses a fixed codebook composed of $2^{K}$ centroids of a Gaussian distribution.
-It then quantizes the posterior sample by finding the closest centroid.
-However, TokenBridge directly quantizes a vanilla Gaussian VAE without limiting the KL divergence of each dimension.
-This leads to suboptimal rate-distortion performance.
-In fact, we can also improve the performance of TokenBridge using TDC.
-Specifically, the quantization centers of TokenBridge are the equal-probability partition centers of the $\mathcal{N}(0,1)$ distribution, which can be seen as a special case of GQ with an evenly distributed codebook $c_{1:K}$.
-From this perspective, the number of PTQ bits should also match the bits-back coding bitrate of the Gaussian VAE, and TDC can therefore enhance the performance of TokenBridge.
-As we demonstrate in Table 3, TDC indeed improves TokenBridge performance by a large margin.
-3.6 Relationship with Reverse Channel Coding
-GQ is closely related to reverse channel coding, which aims to simulate a distribution $q$ using samples from a distribution $p$ (harsha2007communication; li2018strong; havasi2018minimal; Flamich2020CompressingIB; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating).
-For example, Minimal Random Coding (MRC) (havasi2018minimal), when applied to a Gaussian VAE, samples from the categorical distribution with logits given by the likelihood difference:
-$\hat{z}_{i}\sim\hat{q}(c_{1:K})\propto e^{\log q(Z_{i}=c_{j}|X)-\log\mathcal{N}(c_{j}|0,1)}.$ (12)
-The key difference between MRC and GQ is that MRC and its variants (havasi2018minimal; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating) simulate a distribution through stochastic sampling, whereas a VQ-VAE requires deterministic quantization.
-For one dimension quantization, the bias bound of MRC derived from Chatterjee2015TheSS can not be achieved as VQ-VAE does not allow stochastic encoding.
-On the other hand, our achievability and converse bound is compatible with deterministic quantization.
-In terms of quantization error, GQ outperforms MRC by definition (Eq.
-3).
-Besides, GQ without grouping (m=1) can be implemented by bisect search with better asymptotic complexity (See Appendix D.11).
-Additionally, TDC is closely related to the MIRACLE heuristic and IsoKL parametrization of Gaussian VAE (Havasi2018MinimalRC; Flamich2022FastRE; Lin2023MinimalRC).
-More specifically, MIRACLE also proposes adjusting $\lambda$ during VAE training.
-However, MIRACLE only maintains a single $\lambda$, making it less effective for controlling the mean of $D_{KL}$ but less effective for constraining its minimum and maximum values.
-On the other hand, IsoKL imposes strict control on $D_{KL}$ by directly solving for $\sigma$ given $\mu$ using the Lambert W function (corless1996lambert; brezinski1996extrapolation).
-However, IsoKL suffers from numerical issues and leads to suboptimal performance.
-Figure 2: Qualitative results on ImageNet dataset and 0.25 bpp. Our GQ has most visually pleasing reconstruction result.
+TokenBridge (Wang et al., 2025) also converts a pre-trained Gaussian VAE into a VQ-VAE.
+It adopts the Post-Training Quantization (PTQ) technique from model compression, treating the latent as model parameters for discretization.
+TokenBridge uses a fixed codebook of $2^{K}$ centroids from a Gaussian distribution and quantizes the posterior by selecting the closest centroid.
+However, it directly quantizes a vanilla Gaussian VAE without constraining the KL divergence of each dimension, leading to suboptimal rate-distortion performance.
+We can improve TokenBridge using TDC.
+Its quantization centers correspond to equal-probability partition centers of $\mathcal{N}(0,1)$, which is a special case of GQ with an evenly distributed codebook $c_{1:K}$.
+Therefore, the number of PTQ bits should also match $R_{i}$, and TDC can enhance TokenBridge performance (see Table 3).
 4 Experimental Results
 4.1 Experimental Setup
-Models and Baselines For image reconstruction, we select two representative autoencoder architectures: UNet from Stable Diffusion 3 (Esser2024ScalingRF), and ViT from the BSQ (zhao2024image).
-For VQ-VAE baselines, we include vanilla VQGAN (van2017neural) and several representative variants, including FSQ (mentzer2023finite), LFQ (yu2023language), and BSQ (zhao2024image).
-Besides, we compare our approach to pre-trained VQ-VAE of VQGAN-Taming (esser2021taming), VQGAN-SD (rombach2022high) Llama-Gen (sun2024autoregressive), FlowMo (Sargent2025FlowTT), BSQ (zhao2024image) and other conversion approaches such as TokenBridge and ReVQ .
-Additionally, we demonstrate that our TDC technique can improve TokenBridge (wang2025bridging), a previous training-free approach for converting Gaussian VAEs into VQ-VAE.
-For image generation, we employ the Llama transformer (Touvron2023Llama2O; Shi2024ScalableIT).
-Datasets, Bitrates, and Metrics For the datasets, we use the ImageNet (Deng2009ImageNetAL) training split for training, and both the ImageNet and COCO (Lin2014MicrosoftCC) validation split for testing.
-For reconstruction and generation experiments, all images are resized to $256\times 256$ and $128\times 128$ respectively.
-In terms of bitrates, we evaluate the image reconstruction performance of all models using a codebook size of $2^{14}-2^{18}$ and token numbers of $1024$, $2048$, and $4096$, which correspond to bpp (bits-per-pixel) values of 0.221.000.22-1.00.
-This extends the BSQ evaluation beyond 0.250.500.25-0.50 bpp.
-For metrics, we adopt Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang2018TheUE), Structural Similarity Index Measure (SSIM) (Wang2004ImageQA), and reconstruction Fréchet Inception Distance (rFID) (Heusel2017GANsTB) for image reconstruction; generation Fréchet Inception Distance (gFID) and Inception Score (IS) (Salimans2016ImprovedTF) are used for image generation.
-For further details, see Appendix C.
-Table 1: Quantitive results on ImageNet dataset. Our GQ outperforms other VQ-VAEs on both UNet and ViT architecture, across 0.25-1.00 bpp. Bold: best.
-Method
-bpp
-(# of tokens)
-UNet based ViT based
-[…]
-VQGAN
-0.25
-[…]
-26.51 0.125 0.748 5.714 25.39 0.103 0.740 3.518
-FSQ 26.34 0.075 0.756 1.125 25.87 0.109 0.751 3.856
-LFQ 24.74 0.164 0.722 16.337 24.81 0.143 0.725 15.716
-BSQ 25.62 0.086 0.754 1.080 26.52 0.083 0.793 1.649
-GQ (Ours) 27.61 0.059 0.807 0.529 27.88 0.061 0.823 0.932
-VQGAN
-0.50
-($^{\textup{16}}\times2$048)
-29.21 0.052 0.831 1.600 27.86 0.062 0.823 1.228
-FSQ 29.29 0.047 0.845 0.871 28.83 0.055 0.842 1.067
-LFQ 26.90 0.107 0.800 8.035 27.87 0.068 0.829 2.444
-BSQ 27.88 0.059 0.836 0.788 28.44 0.051 0.852 0.700
-GQ (Ours) 30.17 0.039 0.875 0.492 30.42 0.037 0.882 0.592
-VQGAN
-1.00
-[…]
-32.06 0.026 0.896 0.580 31.32 0.032 0.899 0.716
-FSQ 32.38 0.025 0.905 0.636 31.58 0.026 0.905 0.544
-LFQ 28.31 0.074 0.840 3.617 26.67 0.105 0.790 8.288
-BSQ 30.50 0.032 0.900 0.346 31.60 0.027 0.914 0.379
-GQ (Ours) 32.47 0.023 0.907 0.322 31.71 0.024 0.903 0.349
-Table 2: Quantitive results on ImageNet dataset. Our GQ outperforms previous pre-trained VQ-VAEs with less training. Bold: best, : from paper, -: not available.
-Method
-bpp
-(# of tokens)
-[…]
-ImageNet
+Models and Baselines For image reconstruction, we use two representative autoencoder architectures: UNet from Stable Diffusion 3 (Esser et al., 2024), and ViT from BSQ (Zhao et al., 2024).
+For VQ baselines, we train vanilla VQGAN (Esser et al., 2021) and several variants, including FSQ (Mentzer et al., 2023), LFQ (Yu et al., 2023), and BSQ (Zhao et al., 2024) on the same model architecture.
+Additionally, we compare to pre-trained VQ-VAEs such as VQGAN-Taming (Esser et al., 2021), VQGAN-SD (Rombach et al., 2022), Llama-Gen (Sun et al., 2024a), FlowMo (Sargent et al., 2025), BSQ (Zhao et al., 2024), OpenMagViT-V2 (Luo et al., 2024), and conversion methods like TokenBridge and ReVQ.
+We further show that our TDC technique improves TokenBridge (Wang et al., 2025), a previous training-free approach for converting Gaussian VAEs into VQ-VAEs.
+For image generation, we employ the Llama transformer (Touvron et al., 2023; Shi et al., 2024a).
+Datasets, Bitrates, and Metrics We use the ImageNet (Deng et al., 2009) training split for training, and the ImageNet and COCO (Lin et al., 2014) validation splits for testing.
+For reconstruction and generation experiments, images are resized to $256\times 256$ and $128\times 128$, respectively.
+For bitrates, we evaluate reconstruction performance using codebook sizes K from $2^{14}$ to $2^{18}$ and number of token N of $256$, $1024$, $2048$, and $4096$.
+We compute bits-per-pixel (bpp) as $\log_{2}K\times N/H\times W$, where $H,W$ are image height and width.
+In our case, we cover bpp values of 0.070.071.001.00.
+We fix channel dimension to 16 following original Stable Diffusion 3 UNet.
+For bpp $0.50,1.00$ we adopt codebook dimension $4,8$ respectively.
+For bpp <0.50<0.50, we adopt codebook dimension 16.
+For either UNet or ViT architecture, we follow ReVQ (Zhang et al., 2025) to put codebook dimension on channel dimension.
+Metrics include Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Structural Similarity Index Measure (SSIM) (Wang et al., 2004), and reconstruction Fréchet Inception Distance (rFID) (Heusel et al., 2017) for image reconstruction; and generation Fréchet Inception Distance (gFID) and Inception Score (IS) (Salimans et al., 2016) for image generation.
+Further details are provided in Appendix C.
+Table 2: Quantitative results on the ImageNet dataset, compared with pre-trained VQ-VAEs.
+Our GQ outperforms prior pre-trained VQ-VAEs while requiring less training.
+Bold: best, : from paper, -: not available.
 […]
 […]
-VQGAN-Taming
-0.22
+OpenMagViT-V2 (Luo et al., 2024) 0.07 ($^{\textup{14}}\times2$56) 21.63 0.111 0.640 1.17 300 170
+ReVQ-256T (Zhang et al., 2025) 21.96 0.121 0.640 2.05 (Mixed dataset as DC-AE) 83
+GQ (Ours) 22.30 0.116 0.642 1.04 40 87
 […]
-23.38 - - 1.190 (OpenImages) 67
-VQGAN-SD - - - 1.140 (OpenImages) 83
-Llama-Gen-32 24.44 0.064 0.768 0.590 40 70
-FlowMo-Hi 24.93 0.073 0.785 0.560 300 945
-GQ (Ours) 25.31 0.064 0.762 0.491 40 82
-BSQ
-0.28
+VQGAN-SD (Rombach et al., 2022) - - - 1.140 (OpenImages) 83
+Llama-Gen-32 (Sun et al., 2024a) 24.44 0.064 0.768 0.590 40 70
+FlowMo-Hi (Sargent et al., 2025) 24.93 0.073 0.785 0.560 300 945
+GQ (Ours) 25.31 0.064 0.762 0.491 40 87
 […]
-27.78 0.063 0.817 0.990 200 175
 GQ (Ours) 27.86 0.054 0.804 0.424 40 87
-Table 3: Quantitive results of improving TokenBridge based on UNet architecture. Our TDC improves TokenBridge significantly.
-Method
-bpp
-(# of tokens)
-ImageNet validation COCO validation
+Table 3: Quantitative results of improving TokenBridge on the UNet architecture. Applying TDC significantly enhances the reconstruction quality of TokenBridge.
+Method bpp ($K\times N$) ImageNet validation COCO validation
 […]
-Gaussian VAE
-$\approx$ 1.00 (-)
-32.73 0.022 0.910 0.490 32.64 0.018 0.917 2.380
+Gaussian VAE $\approx$ 1.00 (-) 32.73 0.022 0.910 0.490 32.64 0.018 0.917 2.380
 Gaussian VAE (w/ TDC) 32.61 0.023 0.906 0.460 32.69 0.019 0.919 2.717
-TokenBridge
-1.00
-($2^{16}\times 4096$)
-28.24 0.045 0.869 0.823 28.19 0.043 0.878 4.167
+TokenBridge 1.00 ($2^{16}\times 4096$) 28.24 0.045 0.869 0.823 28.19 0.043 0.878 4.167
 TokenBridge (w/ TDC) 31.67 0.025 0.903 0.385 31.56 0.022 0.910 2.171
 GQ (Ours) 32.60 0.022 0.908 0.280 32.53 0.020 0.917 2.153
-4.2 Main Results
-Image Reconstruction In Table 1 and Table 15, we compare our GQ method to other quantization approaches across the 0.250.251.001.00 bpp range.
-The results show that, in terms of reconstruction metrics such as PSNR, LPIPS, SSIM, and rFID, our GQ approach achieves state-of-the-art performance in most cases.
-The advantage of GQ is consistent across both UNet and ViT model architectures, as well as for both the ImageNet and COCO datasets.
-Visually, in Figure 2, it is shown that our GQ also produces pleasing reconstruction and preserves a lot more details in the source image.
-Besides, in Table.
-2, we show that our GQ achieves competitive performance compared with several pre-trained models such as FlowMo, with less training epochs.
-Additionally, in Table.
-11, we show that our GQ outperforms previous methods to discretize Gaussian VAE, including TokenBridge and ReVQ.
-Improving TokenBridge In Table 3, we compare TokenBridge (wang2025bridging) applied to a vanilla Gaussian VAE and a TDC constrainted Gaussian VAE.
-The results show that the quantization error of TokenBridge is quite large when applied to a vanilla Gaussian VAE.
-In contrast, TDC significantly reduces the quantization error.
-Image Generation In Table 4, we evaluate the performance of GQ in terms of image generation.
-It is shown that compared with VQGAN, FSQ, LFQ and BSQ, our GQ has higher codebook usage and codebook entropy.
-In terms of generation FID and IS, our GQ is comparable to FSQ and better than other methods.
-Additionally, we train a DiT (Peebles2022ScalableDM) with same model architecture and same training setting, using the Gaussian VAE with and without TDC.
-It is shown that for limited computation, auto-regressive generation is more efficient than diffusion generation, in terms of both FID and IS.
-This result shows that the conversion from Gaussian VAE to VQ-VAE facilities auto-regressive generation, which improves the efficiency of image generation.
-Complexity Compared with Gaussian VAE, the overhead of GQ is negligible (See Appendix D.10).
-Table 4: Quantitive results on class conditional image generation of ImageNet dataset. Our GQ achieves best codebook usage and competitive generation performance. Bold: best.
+Table 4: Comparison between GQ conversion and its alternatives on the ImageNet dataset.
+GQ achieves better reconstruction quality and can be implemented asymptotically faster when the codebook dimension $m=1$.
+Here, $D_{\infty}(\cdot||\cdot)$ denotes the Rényi $\infty$-divergence.
 […]
-Diffusion
-Gaussian VAE w/o TDC - - 8.35 202.19
-Gaussian VAE w/ TDC - - 8.47 205.94
-Auto-regressive
-VQGAN 16.4% 4.36 8.01 151.40
-FSQ 94.3% 14.74 7.33 224.88
-LFQ 24.9% 9.65 7.73 142.09
-BSQ 99.8% 14.93 7.82 221.64
-TokenBridge 94.6% 14.94 7.82 198.24
-GQ (Ours) 100.0% 15.17 7.67 230.79
-4.3 Ablation Studies
-Effectiveness of Pre-trained Gaussian VAE It is possible to train a vanilla VQ-VAE (van2017neural) using the same codebook as GQ, which is equivalent to a VQ-VAE with a fixed, Gaussian noise codebook.
-It is also possible to directly train the Gaussian VAE neural network with the GQ target in Eq.
-11 using Gumbel-Softmax (jang2016categorical; maddison2016concrete).
-However, as shown in Table 5, both methods do not converge well.
-Furthermore, fine-tuning GQ after initializing it with a pre-trained Gaussian VAE also has only marginal effect on performance.
-These results indicate that GQ’s conversion from a pre-trained Gaussian VAE is necessary and sufficient.
-Table 5: The effect of pre-trained Gaussian VAE. Converting GQ from pre-trained Gaussian VAE is better than training GQ from scratch.
+Gaussian VAE (w/ TDC) O(1)O(1)/$O(1)$ 32.75 0.023 0.906 0.460
+MRC (original) $O(2^{R_{i}})$/$O(1)$ 32.09 0.023 0.906 0.425
+MRC (A coding) $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$$O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ 32.09 0.023 0.906 0.425
+ORC $O(2^{R_{i}})$/$O(1)$ 32.09 0.023 0.906 0.419
+GQ (Ours) $O(R_{i})$/$O(1)$ 32.11 0.023 0.907 0.414
+Table 5: Effect of TDC and alternatives. GQ is effective only when applied to a Gaussian VAE trained with the TDC constraint.
 […]
-GQ (from scratch) Vanilla VQ-VAE 1.00 8.50 0.763 0.156 360.597
-GQ (from scratch) Gumbel-softmax Eq. 11 29.65 0.044 0.866 0.928
-GQ (finetune from Gaussian VAE) Gumbel-softmax Eq. 11 32.45 0.022 0.905 0.264
-GQ (convert from Gaussian VAE) no 32.47 0.023 0.907 0.327
-Effectiveness and Alternatives of TDC To demonstrate the necessity of TDC in Eq.
-9, we train a vanilla Gaussian VAE without TDC.
-As shown in Table 6, the mean $D_{KL(2)}$ of the vanilla Gaussian VAE is close to that of the Gaussian VAE with TDC (3.99 vs.
-4.26 bits).
-However, the range of $D_{KL(2)}$ is much wider for the vanilla model (0.26–27.29 vs.
-2.93–5.63 bits).
-Although the reconstruction performance of the two Gaussian VAEs is very similar (PSNR: 32.73 vs.
-32.61 dB, rFID: 0.490 vs.
-0.460), GQ with TDC outperforms GQ without TDC by a large margin (PSNR: 31.25 vs.
-26.43 dB, rFID: 0.372 vs.
-0.978).
-This demonstrates that TDC is necessary.
-Alternatives to TDC are the MIRACLE heuristic (Havasi2018MinimalRC) and IsoKL (Flamich2022FastRE).
-MIRACLE is not that effective in terms of controlling the range is outperformed by TDC (PSNR 29.48 vs.
-32.11 dB).
-On the other hand, IsoKL imposes a stricter constraint by requiring that $D_{KL}$ is exactly the same across all dimensions.
-Iso-KL enforece the constraint well but has inferior performance (PSNR 30.45 vs.
-32.11 dB).
-This is because IsoKL is not numerically stable and it discards the solution with $\sigma^{2}_{i}>1$.
-In Appendix B, we propose a numerically stable version of Mean-KL (Lin2023MinimalRC), which is a IsoKL that supports grouping ($m>1$).
-However, it does not work well for ViT based models.
-Table 6: The effect of adding constraints to Gaussian VAE. GQ is effective only on Gaussian VAE trained with TDC constraint.
-[…]
-Gaussian VAE no 3.99, 0.26-27.29 - 1.00 32.73 0.022 0.910 0.490
-GQ - - 4 1.00 26.43 0.834 0.054 0.978
+Gaussian VAE None 3.99, 0.26-27.29 - 1.00 32.73 0.022 0.910 0.490
+GQ - - 4 1.00 26.43 0.054 0.834 0.978
 Gaussian VAE MIRACLE 4.34,0.91-26.98 - 1.00 32.82 0.023 0.910 0.436
@@
 GQ - - 4 1.00 30.45 0.878 0.030 0.468
-Gaussian VAE TDC (ours) 4.26, 2.93-5.63 - 1.06 32.61 0.023 0.906 0.460
+Gaussian VAE TDC (Ours, m=1) 4.26, 2.93-5.63 - 1.06 32.61 0.023 0.906 0.460
 GQ - - 4 1.00 32.11 0.023 0.906 0.414
-Alternatives to GQ There are several stochastic alternatives to GQ, such as MRC, ORC, and A coding (havasi2018minimal; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating).
-In Table 7, we compare these methods in terms of reconstruction quality.
+4.2 Main Results
+Image Reconstruction Tables 1 and 15 compare GQ with other quantization methods with same model architecture.
+Across reconstruction metrics such as PSNR, LPIPS, SSIM, and rFID, GQ achieves state-of-the-art performance in most cases.
+Its advantage is consistent across UNet and ViT architectures, as well as on the ImageNet and COCO datasets.
+Moreover, the quantization brought by GQ is marginal compared to Gaussian VAE.
+Figure 2 shows that GQ produces visually pleasing reconstructions, preserving substantially more details from the source image.
+Moreover, Table 2 shows that GQ achieves competitive performance relative to several pre-trained models, such as FlowMo, while requiring fewer training epochs and smaller or comparable model size.
+Finally, GQ outperforms prior Gaussian VAE discretization methods, including TokenBridge and ReVQ.
+Improving TokenBridge Table 3 compares TokenBridge (Wang et al., 2025) applied to a vanilla Gaussian VAE versus a TDC-constrained Gaussian VAE.
+The results show that TokenBridge incurs substantial quantization error on a vanilla Gaussian VAE, whereas applying TDC significantly reduces this error.
+Table 6: Quantitative results for class-conditional image generation on the ImageNet dataset.
+GQ achieves the highest codebook utilization while maintaining competitive generation performance.
+Bold: best.
+Underline: second best.
+[…]
+Diffusion
+Gaussian VAE w/o TDC - 8.35 202.19
+Gaussian VAE w/ TDC - 8.47 205.94
+Auto-regressive
+VQGAN 16.4% 8.01 151.40
+FSQ 94.3% 7.33 224.88
+LFQ 24.9% 7.73 142.09
+BSQ 99.8% 7.82 221.64
+TokenBridge 94.6% 7.82 198.24
+GQ (Ours) 100.0% 7.67 230.79
+Table 7: The reconstruction result when $\log K\neq R_{i}$. GQ works best when $\log K$ matches $R_{i}$.
+[…]
+14 14 $2^{14}\times 1024$ 25.31 0.064 0.762 0.491
+18 14 $2^{14}\times 1024$ 25.24 0.068 0.774 0.527
+14 18 $2^{18}\times 1024$ 27.79 0.059 0.808 0.513
+18 18 $2^{18}\times 1024$ 27.86 0.054 0.804 0.424
+Table 8: GQ is robust to random codebook.
+[…]
+42 27.61 0.059 0.807 0.529
+43 27.61 0.059 0.807 0.523
+44 27.62 0.059 0.807 0.526
+Table 9: The necessity of GQ’s two-stage pipeline is demonstrated: converting a pre-trained Gaussian VAE yields better performance than training GQ from scratch, and further fine-tuning provides only marginal improvements.
+[…]
+VQ-VAE from scratch 1.00 8.50 0.763 0.156 360
+Eq. 10 from scratch 29.65 0.044 0.866 0.928
+Convert from Eq. 10 32.47 0.023 0.907 0.327
+Further finetune Eq. 10 32.45 0.022 0.905 0.264
+Table 10: Ablation study on hyperparameters $\alpha$, $\beta$, and codebook dimension $m$. Here, $\alpha^{\mathrm{bits}}=\alpha/\log 2$.
+[…]
+0.5 1.01 1 25.60 0.088 0.702 1.26
+0.5 1.01 4 26.98 0.068 0.793 0.927
+0.5 1.01 16 27.61 0.059 0.807 0.529
+0.1 1.01 16 27.56 0.058 0.812 0.551
+1.0 1.01 16 27.61 0.063 0.811 0.701
+0.5 1.1 16 27.63 0.060 0.809 0.534
+0.5 1.001 16 27.48 0.058 0.804 0.510
+Image Generation Table 6 evaluates GQ for image generation.
+Compared with VQGAN, FSQ, LFQ, and BSQ, GQ achieves higher codebook usage and codebook entropy.
+In terms of generation FID and IS, GQ is comparable to FSQ and outperforms the other methods.
+Additionally, we train a DiT (Peebles and Xie, 2022) with the same architecture and training setup using Gaussian VAEs with and without TDC.
+The results show that, under limited computational resources, autoregressive generation is more efficient than diffusion in both FID and IS.
+This demonstrates that converting a Gaussian VAE to a VQ-VAE facilitates autoregressive generation, improving image generation efficiency.
+Complexity Compared with a Gaussian VAE, the computational overhead of GQ is negligible (see Appendix D.6).
+4.3 Ablation Study
+GQ Conversion: Alternatives Stochastic alternatives to GQ conversion exist, including MRC, ORC, and A coding (Havasi et al., 2018a; Theis and Yosri, 2021; Flamich et al., 2022; He et al., 2024).
+Table 4 compares these methods in terms of reconstruction quality.
 When applied to a TDC-constrained Gaussian VAE, GQ achieves the best PSNR, SSIM, and rFID.
-Besides, when grouping $m=1$, GQ can be implemented using bisect search.
-This means that GQ is asymptotically faster than MRC, ORC and A coding (See Appendix D.11).
-Table 7: Comparison between MRC methods and GQ on ImageNet dataset. GQ has better reconstruction quality and can be implemented asymptotically faster when grouping $m=1$.
-[…]
-Gaussian VAE (w/ TDC) O(1)O(1)/$O(1)$ 32.75 0.023 0.906 0.460
-MRC (original) $O(2^{DKL(q(Z_{i}|X)||\mathcal{N}(0,1))})$/$O(1)$ 32.09 0.023 0.906 0.425
-MRC (A coding) $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$$O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ 32.09 0.023 0.906 0.425
-ORC $O(2^{DKL(q(Z_{i}|X)||\mathcal{N}(0,1))})$/$O(1)$ 32.09 0.023 0.906 0.419
-GQ (Ours) $O(D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)))$/$O(1)$ 32.11 0.023 0.907 0.414
-The TDC Parameters and Grouping Strategies See Appendix D.
+Moreover, for codebook dimension $m=1$, GQ can be implemented via bisection search, making it asymptotically faster (see Appendix D.7).
+GQ Conversion: Selection of Codebook Size Theorem 3.1-3.2 show that setting codebook size $\log K=R_{i}$ is theoretically optimal.
+In Table 7, we show that empirically GQ is most effective when $\log K$ matches $R_{i}$.
+GQ Conversion: Robustness to Random Codebook The codebook $c_{1:K}$ is sampled from a Gaussian distribution and fixed once generated.
+Table 8 shows that GQ’s performance is nearly invariant to different random seeds.
+GQ Training: TDC To evaluate the necessity of TDC in Eq.
+9, we train a vanilla Gaussian VAE without TDC.
+As shown in Table 5, the mean $R_{i}^{\mathrm{bits}}$ of the vanilla VAE is close to that of the TDC-constrained VAE (3.99 vs.
+4.26 bits), but the range of $R_{i}^{\mathrm{bits}}$ is much wider for the vanilla model (0.26–27.29 vs.
+2.93–5.63 bits).
+While the reconstruction performance of the two Gaussian VAEs is similar (PSNR: 32.73 vs.
+32.61 dB, rFID: 0.490 vs.
+0.460), GQ applied to the TDC-constrained VAE significantly outperforms GQ applied to the vanilla VAE (PSNR: 31.25 vs.
+26.43 dB, rFID: 0.372 vs.
+0.978).
+This demonstrates that TDC is essential for effective GQ conversion.
+Alternatives to TDC include the MIRACLE heuristic (Havasi et al., 2018b) and IsoKL (Flamich et al., 2022).
+MIRACLE is less effective at controlling the range and is outperformed by TDC (PSNR 29.48 vs.
+32.11 dB).
+IsoKL imposes a stricter constraint by requiring that $R_{i}$ is identical across all dimensions.
+While IsoKL enforces this constraint effectively, its overall performance is worse (PSNR 30.45 vs.
+32.11 dB) due to numerical instability and the exclusion of solutions with $\sigma_{i}^{2}>1$.
+In Appendix B, we propose a numerically stable version of Mean-KL (Lin et al., 2023), which extends IsoKL to support multi-dimensional codebooks ($m>1$).
+However, it performs poorly for ViT-based models.
+Necessity of GQ’s Two-Stage Pipeline One could train a vanilla VQ-VAE (Van Den Oord et al., 2017) using the same codebook as GQ, effectively creating a VQ-VAE with a fixed Gaussian noise codebook.
+Alternatively, the Gaussian VAE network could be trained directly with the GQ target in Eq.
+10 using Gumbel-Softmax (Jang et al., 2016; Maddison et al., 2016).
+However, as shown in Table 9, both approaches fail to converge reliably.
+Moreover, fine-tuning GQ after initializing with a pre-trained Gaussian VAE yields only marginal performance improvements.
+Hyperparameters To justify our choice of TDC parameters $\alpha=0.5$ and $\beta=1.01$, we perform a grid search with results shown in Table 10.
+The results indicate that the effect of $\alpha$ and $\beta$ on TDC is minimal as long as $\alpha\leq 0.5$.
+Table 10 also shows that a larger codebook dimension $m$ improves reconstruction.
+Accordingly, we set $m$ to the quotient of the total latent dimension and the total number of tokens, which is the maximum feasible value.
+To achieve optimal reconstruction, we use $\omega=2.0$ for bitrates $\leq 0.50$ bpp and $\omega=0.0$ for 1.00 bpp (see Table 14 for details).
 5 Related works
-VQ-VAE (van2017neural) is an autoencoder that compresses images into discrete tokens.
-Due to the discretization, it is impossible to be train directly using gradient descent.
-Various techniques have been proposed to address this, such as commitment loss (van2017neural), expectation maximization (EM) (roy2018theory), the straight-through estimator (STE) (bengio2013estimating), and Gumbel-softmax (jang2016categorical; maddison2016concrete; sonderby2017continuous; Shi2024TamingSV).
-In addition, VQ-VAE is prone to codebook collapse.
-To mitigate this, various methods have been proposed, such as reducing the code dimension (Yu2021VectorquantizedIM; Sun2024AutoregressiveMB), product quantization (Zheng2022MoVQMQ), residual quantization (Lee2022AutoregressiveIG), dynamic quantization (Huang2023TowardsAI), multi-level quantization (Razavi2019GeneratingDH), feature projection (Zhu2024ScalingTC), rotation codebooks (Fifty2024RestructuringVQ) and etc.
-(yu2021vector; chiu2022self; takida2022sq; zhang2023regularized; huh2023straightening; gautam2023soft; goswami2024hypervq).
-More related to our work, some variants of VQ-VAE with fixed codebook emerge, such as FSQ (mentzer2023finite), LFQ (yu2023language), and BSQ (zhao2024image).
-However, training tricks such as the straight-through estimator (STE) is still required.
-Among all previous works, TokenBridge (wang2025bridging) and ReVQ (zhang2025quantizethenrectifyefficientvqvaetraining) are most aligned with our objective.
-TokenBridge and ReVQ also convert a pre-trained Gaussian VAE into a VQ-VAE.
-However, TokenBridge and ReVQ do not constraint the divergence of Gaussian VAE, leading to suboptimal performance.
-Besides, ReVQ requires some training while our approach is training-free.
-Reverse Channel Coding See Section. 3.6.
+Vector-Quantized Variational Autoencoder VQ-VAE (Van Den Oord et al., 2017) is an autoencoder that compresses images into discrete tokens.
+Due to discretization, it cannot be trained directly via gradient descent.
+Various techniques address this, including commitment loss (Van Den Oord et al., 2017), expectation maximization (EM) (Roy et al., 2018), the straight-through estimator (STE) (Bengio et al., 2013), and Gumbel-Softmax (Jang et al., 2016; Maddison et al., 2016; Sønderby et al., 2017; Shi et al., 2024b).
+VQ-VAE is also prone to codebook collapse.
+To mitigate this, methods such as reducing code dimension (Yu et al., 2021b; Sun et al., 2024b), product quantization (Zheng et al., 2022), residual quantization (Lee et al., 2022), dynamic quantization (Huang et al., 2023), multi-level quantization (Razavi et al., 2019), feature projection (Zhu et al., 2024), and rotation codebooks (Fifty et al., 2024) have been proposed (Yu et al., 2021a; Chiu et al., 2022; Takida et al., 2022; Zhang et al., 2023; Huh et al., 2023; Gautam et al., 2023; Goswami et al., 2024).
+More closely related to our work, some VQ-VAE variants have fixed codebooks, including FSQ (Mentzer et al., 2023), LFQ (Yu et al., 2023), and BSQ (Zhao et al., 2024).
+However, training still relies on tricks such as the straight-through estimator (STE).
+On the other hand, TokenBridge (Wang et al., 2025) and ReVQ (Zhang et al., 2025) convert a pre-trained Gaussian VAE into a VQ-VAE, but they do not constrain the divergence of the Gaussian VAE.
+Reverse Channel Coding GQ is closely related to reverse channel coding, which aims to simulate a distribution $q$ using samples from another distribution $p$ (Harsha et al., 2007; Li and El Gamal, 2018; Havasi et al., 2018a; Flamich et al., 2020; Theis and Yosri, 2021; Flamich et al., 2022; He et al., 2024).
+The key difference between MRC and GQ is that MRC and its variants (Havasi et al., 2018a; Theis and Yosri, 2021; Flamich et al., 2022; He et al., 2024) simulate a distribution via stochastic sampling, whereas a VQ-VAE requires deterministic quantization.
+Regarding quantization error, GQ outperforms MRC by construction (Eq.
+3).
+Moreover, GQ with a one-dimensional codebook ($m=1$) can be implemented via bisection search, achieving superior asymptotic complexity (see Appendix D.7).
+Additionally, TDC is closely related to the MIRACLE heuristic and IsoKL parameterization of Gaussian VAEs (Havasi et al., 2018b; Flamich et al., 2022; Lin et al., 2023).
+Specifically, MIRACLE also adjusts $\lambda$ during VAE training, but it maintains a single $\lambda$, making it less effective at controlling the minimum and maximum values of $R_{i}$.
+IsoKL, in contrast, enforces strict control over $R_{i}$ by directly solving for $\sigma$ given $\mu$ using the Lambert $\mathcal{W}$ function (Corless et al., 1996; Brezinski, 1996), but it suffers from numerical instability and yields suboptimal performance.
 6 Conclusion & Discussion
-In this paper, we propose Gaussian Quant (GQ), an extremely simple yet effective technique that converts a pre-trained Gaussian VAE into a VQ-VAE without any additional training.
-Theoretically, we show that when the logarithm of the GQ codebook size exceeds the bits-back coding bitrate of the Gaussian VAE, a small quantization error is achieved.
-In addition, we propose a target divergence constraint (TDC) to implement GQ effectively.
-Empirically, we demonstrate that GQ outperforms previous discrete VAEs, such as VQGAN, FSQ, LFQ, and BSQ (van2017neural; mentzer2023finite; yu2023language; zhao2024image).
-Furthermore, our TDC also improves the performance of TokenBridge (wang2025bridging).
-We limit our evaluation of GQ to the standard SD3.5 UNet (Esser2024ScalingRF) and the BSQ-ViT architecture.
-Additionally, we restrict the bpp range to 0.220.221.001.00, which extends the BSQ’s original range of 0.250.250.500.50 bpp.
-We acknowledge that there are several highly competitive VQ-VAEs that adopt multi-scale or residual architectures (Tian2024VisualAM; Han2024InfinitySB) or study the low bpp regime (bpp $\leq 0.2$) (Yu2024AnII; Sargent2025FlowTT; zhang2025quantizethenrectifyefficientvqvaetraining).
-However, in this paper, we use standard architectures and a typical bpp range to focus on the core aspects of the quantization method.
-Additionally, we focus on achieving a good trade-off between bitrate and reconstruction quality, leaving the complex relationship between reconstruction and generation performance to future works (wang2024image; xiong2025gigatok; hansen2025learnings).
-Ethics Statement
+In this paper, we propose Gaussian Quant (GQ), a simple yet effective method that constructs a VQ-VAE by first training a constrained Gaussian VAE then converting it into a VQ-VAE.
+The conversion process involves generating codebook from pure Gaussian noise and matching the posterior mean.
+Theoretically, we show that when the logarithm of the GQ codebook size exceeds the bits-back coding bitrate of the Gaussian VAE, the resulting quantization error is small.
+To train a constrained Gaussian VAE for effective conversion, we introduce the target divergence constraint (TDC).
+Empirically, GQ outperforms prior discrete VAEs, including VQGAN, FSQ, LFQ, and BSQ (Van Den Oord et al., 2017; Mentzer et al., 2023; Yu et al., 2023; Zhao et al., 2024), and TDC further improves the performance of TokenBridge (Wang et al., 2025).
+We acknowledge that several highly competitive VQ-VAEs employ multi-scale or residual architectures (Tian et al., 2024; Han et al., 2024).
+In this work, however, we use a standard single-scale architecture to focus on the core aspects of the quantization method.
+Furthermore, recent studies show that strong reconstruction does not necessarily imply strong generation, and achieving good generation often requires feature alignment (Wang et al., 2024; Xiong et al., 2025; Hansen-Estruch et al., 2025).
+Here, we focus on reconstruction and leave the exploration of the complex relationship between reconstruction and generation performance to future work.
+Impact Statement
 The approach proposed in this paper focus on reconstruction of existing images with limited bitrate.
@@
 Nevertheless, the GAN module in decoder might has negative effects, including issues related to mis-representation and trustworthiness.
-Reproducibility Statement
-For theoretical results, the proofs for all theorems are presented in Appendix A.
-For the experiments, we use publicly accessible datasets such as ImageNet (Deng2009ImageNetAL).
-Implementation details and hyper-parameters are provided in Appendix C.
-Besides, we include the source code for reproducing the experimental results in the supplementary material.
 Appendix A Proof of Main Results
-Theorem 1.
-Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i},\sigma_{i}$, and assuming $|\mu_{i}|,\sigma_{i}$ ’s product and sum are bounded by $|\mu_{i}\sigma_{j}|\leq c_{1}$, $|\mu_{i}|+|\sigma_{j}|\leq c_{2}$, then the probability of quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays double exponentially to the amount of nats $t$ that codebook bitrate $\log K$ exceeds the bits-back coding bitrate, i.e.,
-$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+t,$
-$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-e^{t}\underbrace{\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}}_{\textup{constant}})}.$ (6)
+Lemma A.1.
+Given $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)),|\mu_{i}|,|\sigma_{i}|$ are upper-bounded. And therefore $c_{1},c_{2}$ in Theorem 3.1-3.2 are also bounded:
+$|\mu_{i}|\leq\sqrt{2R_{i}},$
+$|\sigma_{i}|\leq\sqrt{-\mathcal{W}_{-1}(-e^{-(2R_{i}+1)})},$
+$c_{1}\leq\sqrt{2R_{i}}\sqrt{-\mathcal{W}_{-1}(-e^{-(2R_{i}+1)})},$
+$c_{2}\leq\sqrt{2R_{i}}+\sqrt{-\mathcal{W}_{-1}(-e^{-(2R_{i}+1)})}.$
+Proof.
+First, we have
+$\mu_{i}^{2}+\sigma_{i}^{2}-1-\log\sigma_{i}^{2}=2R_{i}.$ (15)
+Notice that $\sigma_{i}^{2}-\log\sigma_{i}^{2}\geq 0$. To maximize $\sigma_{i}$, we simply set $\mu_{i}=0$. Then we have $\sigma_{i}^{2}-\log\sigma_{i}^{2}=2R_{i}+1$. We can solve
+$\sigma_{i}^{2}=-\mathcal{W}(e^{-(2R_{i}+1)}),$ (16)
+where $\mathcal{W}(.)$ is Lambert $\mathcal{W}$ function. We select the larger solution
+$\sigma^{2}=-\mathcal{W}_{-1}(e^{-(2R_{i}+1)}),$ (17)
+where $\mathcal{W}_{-1}$ is the lower branch of Lambert $\mathcal{W}$ function. To maximize $\mu_{i}$, we simply set $\sigma_{i}=1$. Then we have
+$\mu_{i}^{2}=2R_{i},\mu_{i}=\sqrt{2R_{i}}.$ (18)
+Theorem 3.1.
+Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i}$ and $\sigma_{i}$, respectively.
+We assume $|\mu_{i}\sigma_{i}|\leq c_{1}$ and $|\mu_{i}|+|\sigma_{i}|\leq c_{2}$.
+Given a fixed $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays doubly exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ exceeds the bits-back coding rate.
+That is,
+$\textup{when }\log K=R_{i}+t,$
+$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\leq\exp{(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5})}.$ (19)
 Proof.
@@
 $=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(\mu_{i}^{2}+\sigma_{i}^{2}-\log\sigma^{2}-1.0+1.0)-\mu_{i}\sigma_{i}})}$
-$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-\mu_{i}\sigma_{i}-0.5})}.$ (18)
+$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-\mu_{i}\sigma_{i}-0.5})}.$ (25)
 Notice that as $\mu_{i}-\sigma_{i}\geq 0,\sigma_{i}>0$, we must have $\mu_{i}\sigma_{i}>0$, then
-$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (19)
+$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (26)
 Similarly, we can show similar result for $\mu_{i}+\sigma_{i}\leq 0$. Obviously we have $\phi(\mu_{i}-\sigma_{i})\leq\phi(x^{\prime})$, and we have
@@
 $=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(\mu_{i}^{2}+\sigma_{i}^{2}-\log\sigma^{2}-1.0+1.0)+\mu_{i}\sigma_{i}})}$
-$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+\mu_{i}\sigma_{i}-0.5})}$
-$\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}$ (20)
+$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}+\mu_{i}\sigma_{i}-0.5})}$
+$\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}$ (27)
 Now, consider the case when $\mu_{i}-\sigma_{i}<0<\mu_{i}+\sigma_{i}$, obviously we must have either $\phi(\mu_{i}-\sigma_{i})\leq\phi(x^{\prime})$, or $\phi(\mu_{i}+\sigma_{i})\leq\phi(x^{\prime})$.
@@
 Therefore, for all $\mu_{i},\sigma_{i}$, we have
-$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (21)
+$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (28)
 Taking the value of $K$ in, we have the result
-$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}$
-$=\exp{(-e^{t}\cdot\sqrt{\frac{2}{\pi}}e^{-|\mu_{\max}\sigma_{\max}|-0.5})}$
+$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}$
+$=\exp{(-e^{t}\cdot\sqrt{\frac{2}{\pi}}e^{-|\mu_{\max}\sigma_{\max}|-0.5})}.$
 $=\exp{(-e^{t}\cdot\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5})}.$ (29)
-Theorem 2.
-The probability of quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increase exponentially to the amount of nats $t$ that codebook bitrate $\log K$ lower than the bits-back coding bitrate, i.e.,
-$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-t,$
-$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\geq 1-e^{-t}\underbrace{\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}}_{\textup{constant}}.$ (7)
+Theorem 3.2.
+Following Theorem 3.1, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increases exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ falls below the bits-back coding rate.
+That is,
+$\textup{when }\log K=R_{i}-t,$
+$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}.$ (30)
 Proof.
-Similar to the proof of Theorem. 1, we have
+Similar to the proof of Theorem. 3.1, we have
 $\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})=(1-(\Phi(\mu_{i}+\sigma_{i})-\Phi(\mu_{i}-\sigma_{i})))^{K}.$ (31)
@@
 $=1-K\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(x_{i}^{2}+\sigma_{i}^{2}-\log\sigma_{i}^{2}-1.0)+|\mu_{i}\sigma_{i}|-0.5}$
-$=1-K\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+|\mu_{i}\sigma_{i}|-0.5}$
-$\geq 1-Ke^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (25)
+$=1-K\sqrt{\frac{2}{\pi}}e^{-R_{i}+|\mu_{i}\sigma_{i}|-0.5}$
+$\geq 1-Ke^{-R_{i}}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (33)
 Similar results can be obtained for $\mu_{i}+\sigma_{i}\leq 0$. For the case that $\mu_{i}-\sigma_{i}\leq 0\leq\mu_{i}+\sigma_{i}$, we have
@@
 $=1-K\sqrt{\frac{2}{\pi}}e^{-\frac{1}{2}(\mu_{i}^{2}+\sigma_{i}^{2}-\log\sigma_{i}^{2}-1.0)-0.5+0.5(\mu_{i}^{2}+\sigma_{i}^{2})}$
-$=1-K\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+0.5(\mu_{i}^{2}+\sigma_{i}^{2})+|\mu_{i}\sigma_{i}|-0.5}$
-$=1-Ke^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (26)
-Taking the value of $K=e^{D_{KL}(q(Z_{i}||\mathcal{N}(0,1)))-t}$ , we have the result
-$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})$ $\geq 1-e^{D_{KL}(q(Z_{i}||\mathcal{N}(0,1)))-t-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$
+$=1-K\sqrt{\frac{2}{\pi}}e^{-R_{i}+0.5(\mu_{i}^{2}+\sigma_{i}^{2})+|\mu_{i}\sigma_{i}|-0.5}$
+$=1-Ke^{-R_{i}}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (34)
+Taking the value of $K=e^{R_{i}-t}$ , we have the result
+$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})$ $\geq 1-e^{R_{i}-t-R_{i}}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$
+$\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5(|\mu_{i}|+|\sigma_{i}|)^{2}-0.5}$
 $\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}.$ (35)
-To better illustrate the significance of these bounds, we provide a practical example.
-We evaluate the ImageNet validation dataset using a pre-trained Gaussian VAE and compute that $c_{1}=8.12$ and $c_{2}=1.50$.
-We then visualize the upper bound and lower bound of $\textup{Pr}{|\hat{z}-\mu_{i}|\geq\sigma_{i}}$ in Fig.
-3.
-The results show that when the codebook bitrate exceeds the bits-back coding bitrate by approximately 10 nats, the probability of large quantization error diminishes to zero.
-Conversely, when the codebook bitrate is smaller than the bits-back coding bitrate, the probability of large quantization error increases.
-Figure 3: A visualization of large quantization error lowerbound and upperbound with ImageNet validation dataset.
 Appendix B Stable Mean-KL Parametrization
-We investigate an alternative to TDC, namely the Mean-KL parametrization (Lin2023MinimalRC), which is considered to be easier to train than TDC since it does not require the construction of an empirical $R(\lambda)$ model.
+We investigate an alternative to TDC, namely the Mean-KL parametrization (Lin et al., 2023), which is considered to be easier to train than TDC since it does not require the construction of an empirical $R(\lambda)$ model.
 B.1 Mean-KL Parametrization
-The Mean-KL parametrization (Lin2023MinimalRC) supports grouping with $m>1$.
-Its neural network output consists of two $m$-dimensional tensors, $\hat{\gamma}{i:i+m}$ and $\tau{i:i+m}$, which allocate the $D_{\mathrm{KL}}$ target $K$ across the $m$ dimensions and determine the mean, respectively.
-More specifically, the Mean-KL parametrization determines the mean $\mu_{i:i+m}$ and variance $\sigma^{2}_{i:i+m}$ as follows, where $\mathcal{W}(\cdot)$ denotes the principal branch of the Lambert W function:
+The Mean-KL parametrization (Lin et al., 2023) supports codebook dimension $m>1$.
+Its neural network output consists of two $m$-dimensional tensors, $\hat{\gamma}_{i:i+m}$ and $\tau_{i:i+m}$, which allocate the $R_{i:i+m}$ target $\log K$ across the $m$ dimensions and determine the mean, respectively.
+More specifically, the Mean-KL parametrization determines the mean $\mu_{i:i+m}$ and variance $\sigma^{2}_{i:i+m}$ as follows, where $\mathcal{W}(\cdot)$ denotes the principal branch of the Lambert $\mathcal{W}$ function:
 $\gamma_{i:i+m}=\textrm{Softmax}(\hat{\gamma}_{i:i+m}),$
@@
 Nonetheless, if only UNet-based models are required, Stable Mean-KL can be an effective alternative to TDC, as it does not require an empirical $R(\lambda)$ model and is significantly simpler to train.
-Table 9: Quantitive results on ImageNet validation dataset.
-Method
-bpp
-(# of tokens)
-UNet based ViT based
+Table 12: Quantitative results on ImageNet validation dataset comparing GQ with different constraint.
+Method bpp ($K\times N$) UNet based ViT based
 […]
-GQ (Mean-KL)
-1.00
 […]
-NaN NaN NaN NaN NaN NaN NaN NaN
 GQ (Stable Mean-KL) 32.35 0.023 0.905 0.280 30.80 0.030 0.891 0.556
@@
 We train all VQ-VAEs on the ImageNet validation dataset using $8\times$ H100 GPUs for approximately 24 hours.
-For UNet models, we train each model for 30 epochs using the ADAM (Kingma2014AdamAM) optimizer with a batch size of $128$ and a learning rate of $1\times 10^{-4}$.
+For UNet models, we train each model for 30 epochs using the Adam (Kingma and Ba, 2014) optimizer with a batch size of $128$ and a learning rate of $1\times 10^{-4}$.
 For ViT models, we train each model for 40 epochs using the Adam optimizer with a batch size of $256$ and a learning rate of $4\times 10^{-7}$.
-All VQ-VAEs are trained using the following distortion objective, which corresponds to the classical VQ-GAN (Esser2020TamingTF) objective employed in the Stable Diffusion VAE (rombach2022high).
+All VQ-VAEs are trained using the following distortion objective, which corresponds to the classical VQ-GAN (Esser et al., 2020) objective employed in the Stable Diffusion VAE (Rombach et al., 2022).
 $\Delta(X,g(z))=\mathcal{L}_{MSE}(X,g(z))+w_{1}\mathcal{L}_{LPIPS}(X,g(z))+w_{2}\mathcal{L}_{GAN}(g(z)).$ (38)
 Following the implementation of Stable Diffusion, we set $w_{1}=1.0$ and $w_{2}=0.75$ for UNet models.
-Consistent with the implementation of BSQ (zhao2024image), we set $w_{1}=0.1$ and $w_{2}=0.1$ for ViT models.
+Consistent with the implementation of BSQ (Zhao et al., 2024), we set $w_{1}=0.1$ and $w_{2}=0.1$ for ViT models.
 For the image generation model, we first train all VQ-VAEs using images of size $128\times 128$, following the same settings as described above.
-Subsequently, we train the auto-regressive transformer for image generation using the implementation of IBQ (Shi2024TamingSV) with a Llama-base transformer architecture.
+Subsequently, we train the auto-regressive transformer for image generation using the implementation of IBQ (Shi et al., 2024b) with a Llama-base transformer architecture.
 The transformer has a vocabulary size of $2^{16}$, $16$ layers, $16$ attention heads, and an embedding dimension of $1024$.
 We train the transformer for $100$ epochs using the Adam optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of $512$.
-C.2 Details of Grouping Strategies
-We extend the notation from main text.
-We group $m$ tokens into one large token with codebook size $K$.
-For each quantization output $\hat{z}_{i}$, we denote the corresponding index in the codebook as $\mathcal{I}_{i}$.
-And we have $c_{\mathcal{I}_{i}}=\hat{z}_{i}$.
-Post Quantization (PQ) The Post Quantization (PQ) grouping strategy happens after GQ.
-We first train a Gaussian VAE with codebook size for each dimension $K^{1/m}$.
-Next, we obtain GQ tokens $\mathcal{I}_{1:d}$.
-Then, we group those $d$ tokens into $d/m$ groups with group size of $m$.
-Denote the group index as $g=0,\ldots,d/m-1$, then each group can be denoted as $\{\mathcal{I}_{gm+l}\},l=1,\ldots,m$.
-In that case, we have $\max\{\mathcal{I}_{gm+l}\}\leq K^{1/m}$.
-Then, we can view each $\mathcal{I}_{gm+l}$ as an integer in a $K^{1/m}$-base numerical system.
-Then, aggregating $m$ tokens $\{\mathcal{I}_{gm+l}\}$ into one large token $I^{m}_{g}$ is as easy as concatenating $m$ tokens into a larger integer in a $K$-based numerical system: $\mathcal{I}_{g}^{m}=\sum_{l=1}^{m}\mathcal{I}_{gm+l}K^{(l-i)/m}\leq K$.
-Post Training (PT) The Post Training (PT) grouping strategy happens after the training of Gaussian VAE.
-We still start with a Gaussian VAE with target codebook size for each dimension $K^{1/m}$.
-Next, instead of performing GQ for each dimension, we consider the following $m$ dimension GQ for each group $g$:
-[…]
-In fact, we can view one dimension GQ as a max likelihood:
-$\hat{z}_{i}=\arg\max_{c_{j}\in\{c_{1:K}\}}\log q(Z_{i}=c_{j}|X),\textrm{ where }c_{1:K}\sim\mathcal{N}(0,1).$
-Then we can extend the max likelihood to $m$ dimension, which is equivalent to the basic version of PT in Eq. 32.
-$\hat{z}_{gm:gm+m}=\arg\max_{c_{j}\in\{c_{1:K}\}}\log q(Z_{gm:gm+m}=c_{j}|X),\textrm{ where }c_{1:K}\sim\mathcal{N}(0,I_{m}).$
-When group size $m$ is large, we observe that vanilla PT in Eq.
-32 leads to codebook collapse (See Table.
-13) and a decay in performance.
-Therefore, we include a regularization term weighted by $\omega$:
-[…]
-Training Aware (TA) The Training Aware (TR) grouping strategy happens before the training of Gaussian VAE.
-More specifically, we directly interoperate grouping with the TDC, and introduce the $m$ group TDC as follows:
-$\mathcal{L}_{TDC}^{m}=\sum_{g=0}^{d/m}\lambda_{g}\sum_{l=1}^{m}D_{KL}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))+\Delta(X,g(z)),$
-$\textup{where }\lambda_{g}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textup{ bits},\\ \lambda_{\textrm{mean}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\in[\log_{2}K-0.5,\log_{2}K+0.5]\textup{ bits},\\ \lambda_{\textrm{max}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textup{ bits}.\end{array}\right$.
-(36)
-Besides, the $\lambda$ update heuristic should also consider $m$ dimension as a group:
-$\lambda_{\textrm{min}}=\lambda_{\textrm{min}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ min}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textrm{ else }\lambda_{\textrm{min}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
-$\lambda_{\textrm{mean}}=\lambda_{\textrm{mean}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ mean}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>\log_{2}K\textrm{ else }\lambda_{\textrm{mean}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
-$\lambda_{\textrm{max}}=\lambda_{\textrm{max}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ max}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textrm{ else }\lambda_{\textrm{max}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}.$ (37)
-C.3 Details of Hyper-parameters
+C.2 Details of Hyper-parameters
 Below, we describe the implementation details along with the definition of hyperparameters for each method.
 In Table 13, we list the values of these hyperparameters for different bits-per-pixel (bpp) settings.
-VQGAN (van2017neural) We adopt the factorized codebook VQGAN variant following (Zheng2022MoVQMQ).
+VQGAN (Van Den Oord et al., 2017) We adopt the factorized codebook VQGAN variant following (Zheng et al., 2022).
 For each codebook, we use a codebook size of $K=2^{16}$ and a group dimension of $m=16$.
@@
 Additionally, we use a codebook loss weight of $\lambda=1.0$ and a commitment loss weight of $\zeta=0.25$.
-FSQ (mentzer2023finite) The only parameter of FSQ is the codebook list $l$, which represents the quantization level for each integer value.
+FSQ (Mentzer et al., 2023) The only parameter of FSQ is the codebook list $l$, which represents the quantization level for each integer value.
 We set each unit value to $2^{4}=16$, and populate $l$ with the appropriate number of $16$s according to the desired bitrate.
-LFQ (yu2023language) For LFQ at $0.25$ bpp, we follow the original paper and split a large codebook of size $2^{16}$ into $n=2$ smaller codebooks, each with $K=2^{8}$ and a codebook dimension of $m=8$.
+LFQ (Yu et al., 2023) For LFQ at $0.25$ bpp, we follow the original paper and split a large codebook of size $2^{16}$ into $n=2$ smaller codebooks, each with $K=2^{8}$ and a codebook dimension of $m=8$.
 We use an entropy loss weight of $\lambda=0.1$ and a commitment loss weight of $\zeta=0.025$.
-BSQ (zhao2024image) We fix the size of each BSQ codebook to $2^{1}$, with a group dimension of $m=1$, and vary the number of codebooks $n$ according to the desired bitrate.
+BSQ (Zhao et al., 2024) We fix the size of each BSQ codebook to $2^{1}$, with a group dimension of $m=1$, and vary the number of codebooks $n$ according to the desired bitrate.
 For the entropy penalization parameter, we set $\lambda=0.1$, following the official implementation.
-GQ We use TR grouping with a fixed codebook size of $K=2^{16}$.
-Each group has dimension $m$, and there are $n$ groups in total, such that $m\times n=16$.
-The value of $m$ varies depending on the bitrate.
+GQ We use a fixed codebook size of $K=2^{16}$. We adopt fixed $\alpha=0.5,\beta=1.01$. We adopt $m=16,8,4,\omega=2.0,2.0,0.0$ for bpp=0.25,0.5,1.0=0.25,0.5,1.0 respectively.
 Table 13: Details of Hyper-parameter values.
@@
 1.00 $K=2^{16},n=4,m=4,\omega=0.0$
-Appendix D Additional Quantitive Results
-D.1 Comparison to other Conversion Methods
-In Table.
-11, we compare our GQ to TokenBridge (wang2025bridging) and ReVQ (zhang2025quantizethenrectifyefficientvqvaetraining), which also convert a Gaussian VAE into a VQ-VAE.
-Those results are from their original paper.
-It is shown that GQ has best reconstruction metrics.
-Besides, only GQ has theoretical guarantee.
-Table 11: Quantitive results on ImageNet dataset. Bold: best, : from paper, -: not available.
-Method
-Training
-free
-Theoretical
-guarantee
-bpp
-(# of tokens)
-[…]
-OpenMagViT-V2 No No 0.07 ($2^{18}\times 256$) 21.63 0.111 0.640 1.17
-TokenBridge Yes No 0.375 ($2^{6}\times 4096$) - - - 1.11
-ReVQ-256T No No 0.07 ($2^{18}\times 256$) 21.96 0.121 0.640 2.05
-GQ (Ours) Yes Yes 0.07 ($2^{18}\times 256$) 22.30 0.116 0.642 1.04
-D.2 The TDC Parameters
-In Table 12, we show the effect of different TDC parameters.
-It is shown that $\beta=1.001,1.01,1.1$ does not make a significant different on result.
-However, setting $\alpha>0.5$ is harmful to performance.
-Table 12: Ablation study on TDC parameters.
-[…]
-0.5 1.01 27.61 0.059 0.807 0.529
-0.1 1.01 27.56 0.058 0.812 0.551
-1.0 1.01 27.61 0.063 0.811 0.701
-0.5 1.1 27.63 0.060 0.809 0.534
-0.5 1.001 27.48 0.058 0.804 0.510
-D.3 Effectiveness of Grouping Strategies
-In Table 13, we evaluate the effect of token grouping techniques.
-The scenario we consider involves grouping four $4$-bit tokens into a single $16$-bit token, which is a reasonable setting for autoregressive generation.
-The results show that PQ has no effect on reconstruction performance, while PT provides some improvements in PSNR and SSIM.
-In contrast, TR, which involves training the Gaussian VAE with a grouping target, achieves the best reconstruction performance.
-Table 13: Effects of token grouping on ImageNet dataset. TR strategy has best reconstruction performance for grouping $m=4$.
-[…]
-Gaussian VAE (w/ TDC) no (m=1) - 4.26, 2.93-5.63 1.06 32.61 0.023 0.906 0.460
-GQ no (m=1) 4 - 1.00 32.11 0.023 0.906 0.414
-GQ PQ (m=4) 16 - 1.00 32.11 0.023 0.906 0.414
-GQ PT (m=4) 16 - 1.00 32.15 0.023 0.907 0.428
-Gaussian VAE (w/ TDC) TR (m=4) - 15.99, 14.81-17.54 1.00 32.62 0.023 0.909 0.331
-GQ TR (m=4) 16 - 1.00 32.47 0.023 0.907 0.322
-Additionally, in Table 14, we show the effect of the regularization parameter $\omega$ for PT and TR.
+Appendix D Additional Quantitative Results
+D.1 Detail in Codebook Usage Hyper-parameter $\omega$
+Additionally, in Table 14, we show the effect of the codebook usage regularization parameter $\omega$.
 For high bitrates, such as $1.00$ bpp, regularization is not required; in other words, setting $\omega=0.0$ yields the good enough codebook usage and rFID.
@@
 $4.0$ 100.0% 15.07 32.44 0.024 0.907 0.343
-Table 15: Quantitive results on COCO 2017 dataset. Bold: best.
-Method
-bpp
-(# of tokens)
-UNet based ViT based
+Table 15: Quantitative results on COCO 2017 dataset. Bold: best.
+Method bpp ($K\times N$) UNet based ViT based
 […]
-VQGAN
-0.25
 […]
-26.25 0.099 0.756 14.110 25.11 0.106 0.747 11.231
 FSQ 26.01 0.072 0.767 5.451 25.85 0.112 0.765 11.213
@@
 GQ (Ours) 27.29 0.057 0.816 3.797 27.55 0.060 0.830 5.305
-VQGAN
-0.50
-($^{\textup{16}}\times2$048)
-29.06 0.049 0.839 6.616 27.83 0.058 0.832 5.461
+VQGAN 0.50 ($^{\textup{16}}\times2$048) 29.06 0.049 0.839 6.616 27.83 0.058 0.832 5.461
 FSQ 29.08 0.043 0.855 4.008 28.51 0.053 0.851 5.390
@@
 GQ (Ours) 30.14 0.037 0.877 3.116 30.18 0.034 0.887 3.616
-VQGAN
-1.00
 […]
-31.97 0.024 0.901 3.455 31.07 0.029 0.904 3.494
 FSQ 32.30 0.022 0.917 2.797 31.48 0.023 0.911 3.045
@@
 $\hat{z}_{i}$ (GQ) $\log_{2}K=4$ bits 32.11 0.023 0.906 0.414
-D.4 The Quantization Error in Pixel Space
+D.2 The Quantization Error in Pixel Space
 Previously we examine the quantization error in latent space. We can further discuss the quantization error in pixel space given the decoder is smooth. More specifically, we have:
-Corollary 3. Following the setting in Theorem 1, and assuming the decoder $g(.)$ satisfy $|g(x_{1})-g(x_{2})|\leq c_{3}|x_{1}-x_{2}|$, we have:
-$Pr\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq\exp(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}).$ (38)
+Corollary 3. Following the setting in Theorem 3.1, and assuming the decoder $g(.)$ satisfy $|g(x_{1})-g(x_{2})|\leq c_{3}|x_{1}-x_{2}|$, we have:
+$\textup{Pr}\{|g(\hat{z})-g(\mu)|\geq c_{3}\sigma_{i}\}\leq\exp(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}).$ (39)
 Proof.
-As $|g(\hat{z}_{i})-g(\mu_{i})|\leq c_{3}|\hat{z}_{i}-\mu_{i}|$, we have $Pr\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq Pr\{c_{3}|\hat{z_{i}}-\mu_{i}|\geq c_{3}\sigma_{i}\}=Pr\{|\hat{z_{i}}-\mu_{i}|\geq\sigma_{i}\}$.
+As $|g(\hat{z}_{i})-g(\mu_{i})|\leq c_{3}|\hat{z}_{i}-\mu_{i}|$, we have $\textrm{Pr}\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq\textrm{Pr}\{c_{3}|\hat{z_{i}}-\mu_{i}|\geq c_{3}\sigma_{i}\}=\textrm{Pr}\{|\hat{z_{i}}-\mu_{i}|\geq\sigma_{i}\}$.
 We can see that theoretically, the quantization error can be magnified by the Lipschitz constant $c_{3}$.
@@
 As shown in the Table 16, the actual loss of quality caused by GQ remains reasonable.
-D.5 Robustness of Codebook to Random Seed
-As the codebook is usually large enough ($2^{16}$) to compensate for the randomness.
-We provide additional in Table 17 showing that the random seed has little effect on reconstruction performance.
-We used three continuos random seeds without cherry-picking.
-It is clear that the performance of GQ is not affected by randomness.
-Table 17: The effect of codebook randomness on the performance of GQ.
-[…]
-42 27.61 0.059 0.807 0.529
-43 27.61 0.059 0.807 0.523
-44 27.62 0.059 0.807 0.526
-D.6 The Effect of Simply Increasing Codebook Size
-According to Theorem 1, increasing $\log K$ too much over $D_{KL}$ will not significantly improve reconstruction but will lead to a waste of bitrate (tokens).
-We provide an additional in Table 18, showing that GQ trained with $14$ bits and quantized into $18$ bits does not perform as well as GQ trained with $18$ bits and quantized into $18$ bits.
-The drawback of simply increasing the codebook size for GQ is that it is not as effective as increasing the target of TDC to that size and quantizing with a proper $\log K=D_{KL}$.
-Table 18: The effect of simply increase codebook size.
-[…]
-$D_{KL}(2)=14$ $\log_{2}K=14$ $2^{14}\times 1024$ 25.31 0.064 0.762 0.491
-$D_{KL}(2)=14$ $\log_{2}K=18$ $2^{18}\times 1024$ 27.79 0.059 0.808 0.513
-$D_{KL}(2)=18$ $\log_{2}K=18$ $2^{18}\times 1024$ 27.86 0.054 0.804 0.424
-D.7 Quantized Latent Visualization
-In Figure 4, we show the t-NSE (Maaten2008VisualizingDU) visualization of latent after GQ, using 5 subclass of ImageNet.
+D.3 Quantized Latent Visualization
+In Figure 3, we show the t-NSE (van der Maaten and Hinton, 2008) visualization of latent after GQ, using 5 subclass of ImageNet.
 Figure 3: The t-NSE visualization of latent of GQ vs. unquantized Gaussian VAE. It is shown that the latent before and after quantization are quite similar.
-D.8 More Generation Results
+D.4 More Generation Results
 To better understand the generation performance, in Table 17, we present the auto-regressive generation result of GQ in different bitrate regime.
@@
 GQ (Ours) $0.25(2^{16}\times 256)$ 5.09 -
-D.9 Prior-Posterior Mismatch
+D.5 Prior-Posterior Mismatch
 Sometimes the Gaussian VAE might suffer from prior-posterior mismatch.
@@
 More specifically, we have
-$D_{KL}(q(Z)||N(0,1))\approx\frac{1}{N}\sum_{i=1}^{N}(\log q(z^{i})-\log N(z^{i}|0,I)).$ (39)
+$D_{KL}(q(Z)||\mathcal{N}(0,1))\approx\frac{1}{N}\sum_{i=1}^{N}(\log q(z^{i})-\log\mathcal{N}(z^{i}|0,I)).$ (40)
 Additionally, we can estimate the optimal bitrate without effected by prior posterior mismatch with a similar approximation:
 $D_{KL}(q(Z|X)||p(Z))\approx\frac{1}{N}\sum_{i=1}^{N}(\log q(z^{i}|X)-\log q(z^{i})).$ (41)
-We train a diffusion model to estimate $\log q(z^{i})$ by using PF-ODE and Skilling-Hutchinson trace estimator (See Appendix D.2 of Song2020ScoreBasedGM).
+We train a diffusion model to estimate $\log q(z^{i})$ by using PF-ODE and Skilling-Hutchinson trace estimator (See Appendix D.2 of Song et al.
+(2020)).
 We use the Gaussian VAE (w/ TDC) + DiT diffusion model and ImageNet validation dataset.
@@
 The result is shown in Table 19.
-The results show that the prior-posterior mismatch is only 0.00033 bits per pixel, accounting for approximately 0.1% of the total bpp.
+The results show that the prior-posterior mismatch is only 0.00033 bits-per-pixel, accounting for approximately 0.1% of the total bpp.
 Furthermore, the best bpp and the actual bpp show no significant difference on a scale of 0.01.
@@
 Table 19: The bitrate and prior-posterior mismatch.
-Divergence bits per pixel
+Divergence bits-per-pixel
 bpp w.r.t. $\mathcal{N}(0,I)$ ($D_{KL}(q(Z\|X)\|\|\mathcal{N}(0,1))$) 0.25
 bpp w.r.t. $q(Z)$ ($D_{KL}(q(Z\|X)\|\|q(Z))$) 0.25
-prior posterior mismatch ($D_{KL}(q(Z)\|\|N(0,1))$) 0.000328
-D.10 Complexity
+prior posterior mismatch ($D_{KL}(q(Z)\|\|\mathcal{N}(0,1))$) 0.000328
+D.6 Complexity
 Table 20: The encoding and decoding overhead of GQ over Gaussian VAE.
@@
 GQ (CUDA) 79 61
-As with FSQ and BSQ (mentzer2023finite; zhao2024image), our codebook can be generated on the fly by maintaining the same random number generator seed on both the encoder and decoder sides.
+As with FSQ and BSQ (Mentzer et al., 2023; Zhao et al., 2024), our codebook can be generated on the fly by maintaining the same random number generator seed on both the encoder and decoder sides.
 Therefore, our GQ model has the same parameter size as the vanilla Gaussian VAE.
@@
 To reduce the computational complexity of GQ, we implement GQ using a tailored CUDA kernel.
-Specifically, we follow the approach of Vonderfecht2025LossyCW, with a key difference: we maintain the codebook, as our bottleneck is not codebook instantiation.
+Specifically, we follow the approach of Vonderfecht and Liu (2025), with a key difference: we maintain the codebook, as our bottleneck is not codebook instantiation.
 Additionally, we avoid the creation of large buffer vectors by performing the summation over $m$ within the CUDA kernel instead of in PyTorch.
@@
 A detailed comparison between the PyTorch implementation and the CUDA implementation of GQ is provided below as GQ_torch and GQ_CUDA, respectively.
-1def GQ_torch(mu, sigma, codebook, m, bs, K):
-2 # mu.shape = (bs, m)
-3 # sigma.shape = (bs, m)
-4 # codebook.shape = (K, m)
-5
-6 # This step create (bs, m, K) vector, which is the performance bottleneck
-7 dist_m =((mu[:,None] - codebook[None])/sigma[:,None])**2
-8 dist = torch.sum(dist_m, dim=1) # sum over m dimension
-9 indices = torch.argmin(dist, dim=1) # argmax over K dimension
-10 zhat = torch.index_select(codebook, 0, indices) # select quantized results
-11 return indices, zhat
-1def GQ_CUDA(mu, sigma, codebook, m, bs, K):
-2 dist = torch.zeros([bs, K])
-3 # need an extension wrapping and register the kernel into operator, we omit it in paper
-4 # see code appendix for details
-5 GQ_Kernel<<<bs * K / 256,256>>>(mu, sigma, codebook, dist, m, bs, K)
-6
-7 indices = torch.argmin(dist, dim=1) # argmax over K dimension
-8 zhat = torch.index_select(codebook, 0, indices) # select quantized results
-9 return indices, zhat
-10
-11__global__ void GQ_Kernel(
-12 const float* mu,
-13 const float* sigma,
-14 const float* codebook,
-15 float* dist,
-16 int64_t m,
-17 int64_t bs,
-18 int64_t K
-19) {
-20 int idx = blockIdx.x * blockDim.x + threadIdx.x;
-21 if (idx >= K * bs) return;
-22 int bi = idx / K;
-23 int ni = idx % K;
-24 float a = 0.0f;
-25 for (int i = 0; i < m; i++) {
-26 float b = (codebook[ni * m + i] - mu[bi * dim + i]) / sigma[bi * dim + i];
-27 a += b * b;
-28 }
-29 dist[idx] = a;
-30 return;
-31}
-D.11 Asymptotic Complexity
-It is noteworthy that GQ without grouping quantization, such as PT or TR, or GQ with a group size of $m=1$, is asymptotically faster than reverse channel coding methods.
+def GQ_torch(mu, sigma, codebook, m, bs, K):
+# mu.shape = (bs, m)
+# sigma.shape = (bs, m)
+# codebook.shape = (K, m)
+# This step create (bs, m, K) vector, which is the performance bottleneck
+dist_m =((mu[:,None] - codebook[None])/sigma[:,None])**2
+dist = torch.sum(dist_m, dim=1) # sum over m dimension
+indices = torch.argmin(dist, dim=1) # argmax over K dimension
+zhat = torch.index_select(codebook, 0, indices) # select quantized results
+return indices, zhat
+def GQ_CUDA(mu, sigma, codebook, m, bs, K):
+dist = torch.zeros([bs, K])
+# need an extension wrapping and register the kernel into operator, we omit it in paper
+# see code appendix for details
+GQ_Kernel<<<bs * K / 256,256>>>(mu, sigma, codebook, dist, m, bs, K)
+indices = torch.argmin(dist, dim=1) # argmax over K dimension
+zhat = torch.index_select(codebook, 0, indices) # select quantized results
+return indices, zhat
+__global__ void GQ_Kernel(
+const float* mu,
+const float* sigma,
+const float* codebook,
+float* dist,
+int64_t m,
+int64_t bs,
+int64_t K
+int idx = blockIdx.x * blockDim.x + threadIdx.x;
+if (idx >= K * bs) return;
+int bi = idx / K;
+int ni = idx % K;
+float a = 0.0f;
+for (int i = 0; i < m; i++) {
+float b = (codebook[ni * m + i] - mu[bi * dim + i]) / sigma[bi * dim + i];
+a += b * b;
+dist[idx] = a;
+return;
+D.7 Asymptotic Complexity
+It is noteworthy that GQ with codebook dimension $m=1$, is asymptotically faster than reverse channel coding methods.
 This is because, for $m=1$, the GQ target in Eq.3 reduces to a quadratic form.
@@
 The details is shown in Algorithm 2.
-On the other hand, most reverse channel coding methods require $O(2^{D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1))})$ computational complexity (havasi2018minimal; Flamich2020CompressingIB; Theis2021AlgorithmsFT).
-A coding (Flamich2022FastRE) can achieve $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ encoding complexity, albeit at the cost of increased decoding complexity.
+On the other hand, most reverse channel coding methods require $O(2^{D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1))})$ computational complexity (Havasi et al., 2018a; Flamich et al., 2020; Theis and Yosri, 2021).
+A coding (Flamich et al., 2022) can achieve $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ encoding complexity, albeit at the cost of increased decoding complexity.
 However, we note that this complexity advantage is not particularly meaningful in practice.
@@
 In Figure 4, we present additional qualitative results showing that GQ achieves superior visual quality.
-However, we also note that none of the approaches is successful in reconstructing the plate of the residential vehicle.
+However, we also note that none of the approaches is successful in reconstructing the license of the residential vehicle.
 The text content remains challenging for low bitrate VQ-VAEs.
-1input $c_{1:K}(\textrm{sorted, such that }c_{j}\leq c_{j+1}),\mu_{i}$
+Algorithm 1 GQ (Argmax)
+0: Codebook $c_{1:K}$ (sorted, $c_{j}\leq c_{j+1}$), $\mu_{i}$
+0: Quantized value $\hat{z}_{i}$, index $j^{*}$
 […]
-3for $j=1$ to $K$ do
-[…]
-[…]
-6return $c_{j^{*}},j^{*}$
-Algorithm 1 GQ (argmax)
-1input $c_{1:K}(\textrm{sorted, such that }c_{j}\leq c_{j+1}),\mu_{i}$
-2l=1,r=Kl=1,r=K
-3while $l+1<r$ do
-4   $m=(l+r)//2$
-5   if $c_{m}<\mu_{i}$ do
-6    $l=m$
-7   else
-8    $r=m$
-[…]
-10   return $c_{l},l$
-11else
-12   return $c_{r},r$
+2:for $j=1$ to $K$ do
+3:  if $T^{*}\leq\|c_{j}-\mu_{i}\|$ then
+4:   $T^{*}\leftarrow\|c_{j}-\mu_{i}\|$
+5:   $j^{*}\leftarrow j$
+6:  end if
+7:end for
+8:return $c_{j^{*}},j^{*}$
 Algorithm 2 GQ (Bisect)
-Figure 5: Qualitative results on ImageNet dataset and 0.25 bpp. None of those approaches correctly reconstruct the plate number.
+0: Codebook $c_{1:K}$ (sorted, $c_{j}\leq c_{j+1}$), $\mu_{i}$
+0: Quantized value $\hat{z}_{i}$, index $l$ or $r$
+$1l\leftarrow$ 1, $r\leftarrow K$
+2:while $l+1<r$ do
+3:  $m\leftarrow\lfloor(l+r)/2\rfloor$
+4:  if $c_{m}<\mu_{i}$ then
+5:   $l\leftarrow m$
+6:  else
+7:   $r\leftarrow m$
+8:  end if
+9:end while
+10:if $\|c_{l}-\mu_{i}\|<\|c_{r}-\mu_{i}\|$ then
+11:  return $c_{l},l$
+12:else
+13:  return $c_{r},r$
+14:end if
+Figure 4: Qualitative results on ImageNet dataset and 0.25 bpp. None of these approaches correctly reconstructs the license plate.